### v1

# Self-testing in the compiled setting via tilted-CHSH inequalities

Abstract
In a Bell scenario, a classical verifier interacts with two non-communicating provers, producing a correlation. Certain correlations allow the verifier to certify, or self-test, the underlying quantum state and measurements. Notably, the family of tilted-CHSH inequalities has been used to self-test any two-qubit state. Self-tests underpin numerous device-independent protocols, however, a significant drawback of such applications is the non-communicating assumption. To address the non-communication assumption Kalai et al. (STOC ’23) give a procedure which compiles a bipartite Bell scenario into a 2-round interaction between a verifier and a single computationally bounded prover. In this work, we formalize self-testing for compiled Bell scenarios. We prove the maximal quantum value is preserved under compilation for the family of tilted-CHSH inequalities, and that any maximal violation constitutes a compiled self-test. More specifically, we establish the existence of an efficient isometry recovering the state and measurements in the second round.
1 Introduction
In a bipartite Bell scenario, two non-communicating provers receive inputs $x$ and $y$ and reply with outputs $a$ and $b$. The probabilities of $a,b$ given $x,y$ determine a correlation $\mathsf{p}={p(a,b|x,y)}$. Bell’s celebrated theorem implies that if the provers are permitted to share an entangled quantum state and make local quantum measurements, called a quantum model, then some behaviours have no realization under a local hidden-variable (LHV) model [Bel64]. The distinction between quantum and classical behaviours is often analyzed via Bell inequalities, which are linear inequalities on the behaviour satisfied by all LHV models, but can be violated by particular entangled quantum systems. The quintessential example is the CHSH inequality, considered in [CHSH69]. Bell inequality violations play a major role in areas of quantum computing and communication such as device-independent cryptography [ABG+07, Col07, PAM+10, PAB+09, Sca12], protocols for verifiable delegated quantum computation [RUV13, Gri19], and in the study of multiparty interactive proofs (MIPs) and the variant MIP with entangled provers [CHTW04, JNV+21].
Many key applications of Bell inequalities rely on a remarkable property known as self-testing [MY04, MYS12, YN13, ŠB20]. Informally, a Bell inequality is a self-test for an honest (ideal) quantum model $\mathrm{Q}$ if there exist local quantum operations which transform any optimal quantum model $\mathrm{Q}^{\prime}$ into the honest model $\mathrm{Q}$. It is well-known that the CHSH inequality is a self-test for a quantum model containing the maximally entangled state on two qubits [MY04]. Another notable example is the family of tilted-CHSH inequalities which have been used to self-test any bipartite state on two qubits [AMP12, YN13, BP15]. This latter family was instrumental in the celebrated result of Coladengelo, Tong Goh, and Scarani who employed these inequalities to self-test any pure bipartite state [CGS17].
Despite the enormous success of self-testing, a significant drawback in several applications is the assumption of non-communicating provers. Recently, a number of cryptographic approaches have been proposed that replace the non-communication assumption with computational assumptions [KLVY23, MV21, KMCVY22]. One approach, the KLVY compilation procedure introduced in [KLVY23], transforms a $2$-prover $1$-round Bell scenario, into a $1$-prover $2$-round scenario, with a computationally bounded prover. The core ingredient in the KLVY compilation procedure is quantum homomorphic encryption (QHE), which is used to simulate non-communication between the rounds of interaction. Essentially, the QHE scheme allows a quantum prover to perform honest measurements without knowing the given measurement settings.
For the case of classical provers, it was shown that the KLVY compilation procedure preserves the winning probabilities [KLVY23]. In a follow-up work by Natarajan and Zhang [NZ23], it was shown that the maximal quantum violation of the CHSH inequality in the compiled setting is bounded by the maximal violation in the bipartite setting. Subsequent works have studied the KLVY compilation procedure in various contexts, including demonstrating quantum soundness for all 2-player XOR nonlocal games [CMM+24], Bell inequalities tailored to maximally entangled bipartite states [BVB+24], delegated quantum computation with a single-device [NZ23, MNZ24], and even in the study of contextuality [ABCC24]. Despite these advancements, many foundational results in the bipartite setting, such as the self-testing for all bipartite qubit states  [YN13, BP15] have yet to be reproduced in the compiled setting.
Summary of results
We show that for a family of extended tilted-CHSH inequalities [BSB24], the maximal quantum violation in the compiled setting is bounded by the maximal quantum violation in the bipartite setting. To obtain this bound, we establish a technical result which extends the powerful sum-of-squares technique to allow terms with arbitrarily high degrees in Bob’s operators, partially answering a question posed in [CMM+24]. Our next contribution is a step towards a mathematical theory of self-testing in the compiled setting. We introduce a notion of self-testing in the compiled setting, which certifies the second-round measurements and the post-measurement state after the first round of interaction. Lastly, we combine these points by showing that the extended tilted-CHSH inequalities are self-tests according to our definition. In summary:
  • In Lemma 3.2 we extend the SOS proof techniques of Cui et al. [CMM+24] to allow for monomials with arbitrarily high degree in Bob operators.
  • In Theorem 4.2 we prove that the quantum bound for an extended family of tilted-CHSH expressions is preserved under the KLVY compilation procedure.
  • In Definition 3.4 we introduce a notion of self-testing for compiled Bell inequalities, which certifies the structure of the measurements in the second round and the state of the device after the first round of the protocol.
  • In Theorem 4.4 we combine the previous two points to show that the extended tilted-CHSH inequalities are a self-test according to our definition.
1.1 Technical discussion
In this section, we briefly discuss the KLVY compilation procedure and the family of extended tilted-CHSH inequalities. Additionally, we introduce the compiled counterpart to a bipartite quantum model, which is essential for our approach to self-testing. Lastly, we review self-testing in the bipartite setting and give an informal version of our definition of self-testing in the compiled setting.
The KLVY Compiler
The KLVY compiler transforms any Bell scenario into an interaction between a verifier and a single computationally-bounded prover, who sequentially assumes the roles of the two provers Alice and Bob. The procedure is roughly as follows:
  1. 1.
    The prover prepares an initial state $|\psi\rangle$.
  2. 2.
    The verifier samples challenges $(x,y)$, sends the encrypted first challenge $\chi=\mathsf{Enc}(x)$ to the prover, who acts as Alice, measures the state, and returns the outcome $\alpha$.
  3. 3.
    The verifier then sends $y$ in the clear, and the prover, acting as Bob, measures the residual state and returns the outcome $b$.
  4. 4.
    The verifier determines the decrypted answer pair $(a=\mathsf{Dec}(\alpha),b)$.
The compiled value of a Bell inequality is always at least the quantum value, as any quantum strategy can be implemented with homomorphic encryption [KLVY23]. Establishing upper bounds on the compiled value is challenging, since general techniques for bounding the value of a nonlocal game depend on the spatial separation between the two provers.
Extended tilted-CHSH inequalities
The family of “tilted-CHSH” inequalities was introduced by Acín, Massar, and Pironio [AMP12]. The Bell operators for these inequalities take the form $\alpha_{\theta}\langle A_{0}\rangle+\langle A_{0}B_{0}\rangle+\langle A_{0}B_{1}\rangle+\langle A_{1}B_{0}\rangle-\langle A_{1}B_{1}\rangle$, where $\langle A_{x}B_{y}\rangle,\langle A_{x}\rangle$ denote the expectation of measurements corresponding to settings $X=x,Y=y$, and $\alpha_{\theta}\in\mathbb{R}$. This family of inequalities has applications in self-testing, notably, they are tailored to robustly self-test the 2-qubit states $\cos(\theta)|00\rangle+\sin(\theta)|11\rangle$ [YN13, BP15]. Subsequently, these inequalities were used as part of a more complex protocol to obtain self-testing for all pure bipartite quantum states [CGS17]. The work of Barizien, Sekatski, and Bancal [BSB24] extended this family to include extra degrees of freedom in Bob’s measurements, which we will refer to as “extended” tilted-CHSH inequalities.
Compiled counterpart of a bipartite model
A bipartite (quantum) model $\mathrm{Q}$, consists of a shared state $|\phi\rangle$, along with POVM measurements $\{P_{a|x}\}$ and $\{Q_{b|y}\}$ for Alice and Bob, respectively. In the bipartite setting, one can consider the situation where Alice measures first using POVM $\{P_{a|x}\}$, collapsing the state to a post-measurement state $\rho_{a|x}$ on Bob’s space, upon which Bob performs his measurement $\{Q_{b|y}\}$. In the case that $\rho_{a|x}$ is pure, we write $\rho_{a|x}=|\phi_{a|x}\rangle\!\langle\phi_{a|x}|$. Analogously, we consider a compiled (quantum) model $\widetilde{\mathrm{Q}}$ consisting of a family of post-measurement states for “Alice” $|\widetilde{\phi}_{\alpha|\chi}\rangle$, which correspond to the state of the device following the encrypted question $\chi$, and encrypted answer $\alpha$, and a POVM $\{N_{b|y}\}$ employed by “Bob”. One could also consider a more general compiled quantum model which includes a description of the initial state and Alice’s operators. The point of taking the coarser model is that it allows us to introduce the notion of the compiled counterpart of a bipartite model $\mathrm{Q}$, which relates the post-measurement information in bipartite setting with that of the compiled setting.
Definition 1.1 (Compiled counterpart (informal)).
Given a bipartite model $\mathrm{Q}$ for which $\rho_{a|x}$ is pure, the compiled counter-part $\widetilde{\mathrm{Q}}|_{B}$, is a compiled model $\widetilde{\mathrm{Q}}|_{B}=(\{|\widetilde{\phi}_{\alpha|\chi}\rangle\},\{\{N_{b|y}\}_{b\in\mathcal{B}}\}$) satisfying the following conditions:
$|\widetilde{\phi}_{\alpha|\chi}\rangle$ $=|\phi_{a|x}\rangle,\ \text{ for all }\chi:\mathsf{Enc}(x)=\chi,\ \alpha:\mathsf{Dec}(\alpha)=a.$
$N_{b|y}$ $=Q_{b|y},\ \text{ for all }b,y.$
Strictly speaking, Definition 1.1 is not a full counter-part of a quantum model $\mathrm{Q}$, but is more accurately described as a counter-part to Bob’s view of $\mathrm{Q}$. We also note that this definition requires that the states $\rho_{a|x}$, which are obtained by applying Alice’s measurement and tracing out, are pure. This condition will be satisfied when we take $\mathrm{Q}$ to be the honest quantum model for the tilted-CHSH inequalities.
Self-testing in the Bipartite Setting
A Bell inequality, $I\leq\eta^{Q}$, self-tests an optimal, also called honest, bipartite model $\mathrm{Q}^{*}$, if any optimal bipartite model is essentially the same as $\mathrm{Q}^{*}$, modulo, some physically irrelevant degrees of freedom. This is more formally treated by the existence of local isometries which maps the observed model to the honest one.
Definition 1.2 (Bipartite self-test (informal)).
A Bell inequality, $I\leq\eta$, is a self-test for a bipartite model $\mathrm{Q}^{*}=\left(\{P_{a|x}\},\{Q_{b|y}\},|\phi\rangle\right)$ if, for any bipartite model $\mathrm{Q}=\left(\{M_{a|x}\},\{N_{b|y}\},|\Psi\rangle\right)$ that saturates the bound $\eta$, there exist an auxiliary state $|\zeta\rangle$ and local isometries $V_{A}$ and $V_{B}$, that satisfy the following condition for all $x,y,a,b$:
$V_{A}\otimes V_{B}(M_{a|x}\otimes N_{b|y})|\Psi\rangle=(P_{a|x}\otimes Q_{b|y})|\phi\rangle\otimes|\zeta\rangle,$
Self-testing in the Compiled Setting
There are two difficulties in extending the notion of self-testing to the compiled setting. First, what are the relevant or irrelevant degrees of freedom in the compiled setting? Indeed, if we consider arbitrary isometries, then it is always possible to extract the ideal model $\widetilde{\mathrm{Q}}^{*}$. Secondly, using current techniques it is unclear how to make determinations about Alice’s operators. In particular, rigidity relations enjoyed by Alice’s operators in the bipartite model, such as anti-commutation, may not hold for the measurements performed by the corresponding homomorphic evaluation. Below we outline a notion of self-testing that naturally addresses the first concern, by requiring the isometry to be efficient, and somewhat sidesteps the second concern by focusing on the post-measurement perspective after which Alice operators have already been applied. We note that our definition covers self-testing type results from [NZ23, CMM+24, BVB+24].
Definition 1.3 (Compiled self-test (informal)).
Let $\mathrm{Q}^{*}=(\{P_{a|x}\},\{Q_{b|y}\},|\phi\rangle)$, denote the honest bipartite model. A Bell inequality $I\leq\eta$ is a self-test for the corresponding compiled counter-part $\widetilde{\mathrm{Q}}^{*}|_{B}=(\{|\widetilde{\phi}_{\alpha|\chi}\rangle\},\{Q_{b|y}\})$, if for every complied model $\widetilde{\mathrm{Q}}=\left(\{|\Psi_{\alpha|\chi}\rangle\},\{N_{b|y}\}\right)$ that achieves $I=\eta$, there exists a negligible function $\delta_{\mathsf{QHE}}$, an efficient isometry $V$, and auxiliary states $|\mathsf{aux}_{\alpha|\chi}\rangle$, which satisfy the following for all $x,b,y$:
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\big{\|}VN_{b|y}|\Psi_{\alpha|\chi}\rangle-Q_{b|y}|\widetilde{\phi}_{\alpha|\chi}\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big{\|}^{2}\leq O(\delta_{\mathsf{QHE}}).$
1.2 Outline of proof techniques
Upper Bounding the Complied Value
Upper bounds on the compiled quantum value of certain Bell inequality can often be verified using the well-known sum-of-squares (SOS) technique [NZ23, CMM+24, BVB+24]. Informally, this technique [^1] We suggest reading Section 4 of [CMM+24] for a more formal discussion of this technique. relates the maximum compiled value $\eta$, of a Bell functional $I$, to a suitably nice decomposition of the Bell operator or Bell polynomial $S$ as a sum of hermitian squares,
$\eta\mathbbm{1}-S=\sum_{i}P^{\dagger}_{i}P_{i}.$
Here, suitably nice requires that the terms $P_{i}$ involved in the decomposition are degree $2$ monomials in Alice and Bob’s observables.
Our first technical result provides an extension of the aforementioned SOS technique to a larger family of Bell polynomials. More specifically, we extend the pseudo-expectation techniques in [NZ23, CMM+24] to allow for evaluations on polynomial terms $P_{i}$ that consist of arbitrary monomials in the algebra generated by Bob’s observables. In Lemma 3.2 we prove that an extended pseudo-expectation will be positive on the corresponding hermitian square $P_{i}^{\dagger}P_{i}$ for any such term $P_{i}$. As a consequence of Lemma 3.2, we obtain the following general result that can accommodate SOS decompositions of higher degree in Bob’s observables [^2] Although our set-up and proof of Lemma 3.2 are given for binary observables, we expect this approach can generalize to larger question sets..
Theorem 1.4 (Informal).
Let $\mathcal{S}$ be a Bell scenario and let $S$ be a Bell polynomial. Suppose that $\eta\mathbbm{1}-S=\sum P^{\dagger}_{i}P_{i}$ is a SOS decomposition with each term $P_{i}$ of the form
$P_{i}=\sum_{j}\gamma_{j}(A_{x})^{k_{j}}w_{j}(B),$ (1.1)
for some $\gamma_{j}\in\mathbb{C}$, $k_{j}\in\{0,1\}$ and, $w_{j}(B)$, an arbitrary monomial in Bob’s observables. Then $\eta$ is an upper bound on the compiled value of $S$.
Next, we apply our general result to the SOS decomposition for the extended tilted-CHSH inequalities presented in [BSB24]. In Theorem 4.2 we use our new result to prove that the maximum quantum value achieved for any of the extended tilted-CHSH functionals is preserved by the KLVY compilation procedure.
Compiled extended tilted-CHSH inequalities are compiled partial self-tests
In the compiled setting, we show that our definition of self-testing is satisfied by the extended family of tilted-CHSH inequalities. Our proof approach is inspired by the self-testing results of Bamps and Pironio [BP15] which show a near-optimal Bell violation in the compiled setting implies key relations on the expectation of each SOS polynomial. A key difference to the bipartite case is that, instead of proving relations of Alice’s and Bob’s operators on a shared bipartite state. The relations we prove capture the structure of the observables in the second round measured on the post-measurement state after the first round. Interestingly, these relations strongly resemble those found in the bipartite setting, indicating that our notion of a compiled self-test allows existing techniques to be applied.
Future outlook
Moving forward, we consider several natural directions for following up on this work:
  • One central goal in the compiled setting is to determine if the KLVY compiler always preserves the bipartite quantum value. Bell inequalities offer more flexibility than nonlocal games, perhaps this could be utilized to find compiled Bell inequalities with greater quantum violations in the compiled case than in the bipartite case.
  • Tilted-CHSH inequalities were an integral component of the self-testing for all pure bipartite states [CGS17]. Building off of our work on compiled tilted-CHSH inequalities, a natural question is whether similar results can be obtained in the compiled setting.
  • It would be desirable to understand the fundamental limitations of our notion of self-testing and other similar notions such as the computational self-testing given in [MV21]? Furthermore, is a finer notion of self-testing in the compiled setting that characterizes both Alice and Bob’s operators and the initial state possible without specifying the underlying QHE scheme?
  • All current techniques for bounding the value of compiled nonlocal games/Bell inequalities can be obtained using some variant of the sum-of-squares decomposition approach. Given our improvements to this approach outlined in Lemma 3.2, it is possible to search for valid decompositions which include arbitrary words in Bob’s operators. Is it possible to use this approach to give a limited variant of the NPA hierarchy [NPA08] in the compiled setting?
Outline of the paper
The paper proceeds as follows. Section 2 covers our mathematical notation, preliminaries on bipartite Bell inequalities, definitions for efficient quantum operations, and a brief overview of quantum homomorphic encryption. In Section 3, we define compiled Bell inequalities for bounded quantum provers, discuss bounding the quantum value of a compiled Bell inequality, and define a notion of self-testing for compiled quantum models. In Section 4 we consider the family of compiled tilted-CHSH inequalities, and show that maximal Bell inequality violation in the compiled setting is bounded by the maximal violation in the bipartite setting. We also show that this family of compiled Bell inequalities satisfies our definition of compiled self-tests outlined in Section 3.
Acknowledgements
The authors thank Simon Schmidt, Ivan Šupić, Anand Natarajan and Tina Zhang for helpful discussions. AM is supported by NSERC Alliance Consortia Quantum grants, reference number: ALLRP 578455 - 22 and the NSERC Discovery Grants Program. CP acknowledges funding support under the Digital Horizon Europe project FoQaCiA, Foundations of quantum computational advantage, grant no. 101070558, together with the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). LW acknowledges funding support from the EPSRC grant EP/SO23607/1.
2 Preliminaries
2.1 Mathematical notation
Throughout the article, Hilbert spaces are denoted by $\mathcal{H}$, and are assumed to be finite-dimensional unless explicitly stated otherwise. Elements of $\mathcal{H}$ are denoted by $|v\rangle\in\mathcal{H}$, where the inner product $\langle u|v\rangle$ for $|v\rangle,|u\rangle\in\mathcal{H}$ is linear in the second argument and defines the vector norm $\||v\rangle\|=\sqrt{\langle v|v\rangle}$. Quantum states are the norm 1 elements of $\mathcal{H}$. In this work, $\mathbb{B}(\mathcal{H})$ denotes the unital $\dagger$-algebra of bounded linear operators on $\mathcal{H}$ with norm $\|M\|^{2}_{\mathrm{op}}=\sup_{|v\rangle\in\mathcal{H}}\langle v|M^{\dagger}M|v\rangle/\langle v|v\rangle$. The unit in $\mathbb{B}(\mathcal{H})$ is denoted by $\mathbbm{1}$, and we write $|M|=M^{\dagger}M$ for the positive part of $M\in\mathbb{B}(\mathcal{H})$. Given a set $A$, a collection of positive operators $\{M_{a}\geq 0:a\in\mathcal{A}\}$ with the property that $\sum_{a}M_{a}=\mathbbm{1}$, is called a POVM over $\mathcal{A}$. When the operators in a POVM are orthogonal projections, we call it a PVM. Given another finite set $\mathcal{X}$, we let $\mathscr{A}_{POVM}^{(\mathcal{X},\mathcal{A})}$ be the $\dagger$-subalgebra of $\mathbb{B}(\mathcal{H})$ consisting of $\dagger$-polynomials in the POVM elements $\{M_{a|x}\geq 0:a\in\mathcal{A}\}$. Given a random variable $X$ with a distribution $\mu:X\to\mathbb{R}_{\geq 0}$ such that $\sum_{x\in X}\mu(x)=1$ we denote the expectation of $X$ by $\mathop{\mathbb{E}}[X]=\sum_{x\in X}\mu(x)\cdot x$. For $a,b\in\mathbb{R}$ and $\delta>0$, $a\approx_{\delta}b$ is short for $|a-b|\leq\delta$.
2.2 Efficient quantum circuits and homomorphic encryption
To define the KLVY compilation procedure we require the following concepts from quantum cryptography.
Definition 2.1.
A procedure $\mathcal{P}$ is quantum polynomial time (QPT) if:
  1. 1.
    there exists a uniform logspace family of quantum circuits that implement $\mathcal{P}$, and
  2. 2.
    the runtime of the circuit is polynomial in the number of qubits and the security parameter $\lambda\in\mathbb{N}$.
In particular, the following definition is relevant in Section 3.
Definition 2.2.
Let $\mathcal{H}$ and $\mathcal{H}^{\prime}$ be Hilbert spaces. An isometry $V:\mathcal{H}\mapsto\mathcal{H}^{\prime}$ is efficient if there is a QPT procedure implementing $V$.
For more details on efficiently implementable quantum operations we refer the reader to the discussions in [CMM+24]. We also need to define a quantum homomorphic encryption (QHE) scheme. A formal definition of QHE first appeared in [BJ15]. We follow the description of QHE outlined in [KLVY23, Bra18]:
Definition 2.3.
A quantum homomorphic encryption scheme $\mathsf{Q}$ for a family of circuits $\mathcal{C}$ consists of a security parameter $\lambda\in\mathbb{N}$ and the following algorithms:
  1. (i)
    A PPT algorithm $\mathsf{Gen}$ which takes as input a unary encoding $1^{\lambda}$ of the security parameter $\lambda\in\mathbb{N}$ and outputs a secret key $\mathsf{sk}$.
  2. (ii)
    A PPT algorithm $\mathsf{Enc}$ which takes as input the secret key $\mathsf{sk}$ and a plaintext $x\in\{0,1\}^{n}$ and produces a ciphertext $\chi\in\{0,1\}^{k}$.
  3. (iii)
    A QPT algorithm $\mathsf{Eval}$ which takes as input a classical description of a quantum circuit $\mathsf{C}:\mathcal{H}\otimes(\mathbb{C}^{2})^{\otimes n}\to(\mathbb{C}^{2})^{\otimes m}$ from $\mathcal{C}$, a quantum plaintext $|\Psi\rangle\in\mathcal{H}$ on Hilbert space, a ciphertext $\chi$, and evaluates a quantum circuit $\mathsf{Eval}_{\mathsf{C}}(|\Psi\rangle\otimes|0\rangle^{\mathrm{poly}(\lambda,n)},\chi)$ producing a ciphertext $\alpha\in\{0,1\}^{\ell}$.
  4. (iv)
    A QPT algorithm $\mathsf{Dec}$ which takes as input ciphertext $\alpha$, and secret key $\mathsf{sk}$, and produces a quantum state $|\Psi^{\prime}\rangle$.
Although the existence of algorithms (i)-(iv) defines a QHE scheme, we consider several additional important properties a scheme may or may not possess:
  1. 1.
    (Correctness with auxiliary input). For every security parameter $\lambda\in\mathbb{N}$, secret key $\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})$, classical circuit $\mathsf{C}:\mathcal{H}_{A}\otimes(\mathbb{C}^{2})^{\otimes n}\to\{0,1\}^{m}$, quantum state $|\Psi\rangle_{AB}\in\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, plaintext $x\in\{0,1\}^{n}$ ciphertext $\chi\leftarrow\mathsf{Enc}(x,\mathsf{sk})$, the following procedures produce states with negligible trace distance with respect to $\lambda$:
    1. (a)
      Starting from the pair $(x,|\Psi\rangle_{AB})$, run the quantum circuit $\mathsf{C}$ on register $A$, outputting the classical string $a\in\{0,1\}^{m}$ along with the contents of register $B$.
    2. (b)
      Starting from $(\chi,|\Psi\rangle_{AB})$, run the circuit $\mathsf{Eval}_{C}(\cdot)$ on register $A$, obtaining ciphertext $\alpha\in\{0,1\}^{\ell}$, output $a^{\prime}=\mathsf{Dec}(\alpha,\mathsf{sk})$ along with the contents of register $B$.
  2. 2.
    (IND-CPA security). Fix a secret key $\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})$. Any quantum polynomial time adversary $\mathfrak{A}$ with access to $\mathsf{Enc}(\cdot,\mathsf{sk})$ (but does not know $\mathsf{sk}$) cannot distinguish between ciphertexts $\chi\leftarrow\mathsf{Enc}(x_{0},\mathsf{sk})$ and $\chi^{\prime}\leftarrow\mathsf{Enc}(x_{1},\mathsf{sk})$ with non-negligible probability in $\lambda$, where $x_{0}$ and $x_{1}$ are any plaintexts chosen by the adversary. That is
    $|\Pr[\mathfrak{A}^{\mathsf{Enc}(x_{0},\mathsf{sk})}(x_{0})=1]-\Pr[\mathfrak{A}^{\mathsf{Enc}(x_{0},\mathsf{sk})}(x_{1})=1]|\leq\mathrm{negl}(\lambda),$
    for all pairs $(x_{0},x_{1})$.
  3. 3.
    (Compactness). The runtime of $\mathsf{Dec}$ does not depend on the description of $\mathsf{C}$. That is, $|\mathsf{Eval}_{C}(|\Psi\rangle\otimes|0\rangle^{\mathrm{poly}(\lambda,n)},\chi)|=\mathrm{poly}(\lambda)$.
  4. 4.
    (Universality). There is a QHE scheme with (i)-(iv) for any quantum circuit.
For the KLVY compilation procedure, we only require schemes that satisfy (1) and (2). A QHE satisfying (3) and (4) is called a quantum fully homomorphic encryption (QFHE) scheme. Schemes satisfying (1)-(4) have been proposed in [Mah20, Bra18].
2.3 Bell scenarios, inequalities, and violations
Let $\mathcal{A},\mathcal{B},\mathcal{X},$ and $\mathcal{Y}$ be finite sets, with $|\mathcal{A}|=m_{A}$, $|\mathcal{B}|=m_{B}$, $|\mathcal{X}|=n_{A}$, and $|\mathcal{Y}|=n_{B}$. A bipartite Bell scenario is described by the tuple $\mathcal{S}=(\mathcal{A},\mathcal{B},\mathcal{X},\mathcal{Y},\pi)$, where $\pi:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is a distribution over the measurement settings. In this scenario, each party receives an input $x\in\mathcal{X}$ (resp. $y\in\mathcal{Y}$) sampled according to $\pi$, and returns outputs $a\in\mathcal{A}$ (resp. $b\in\mathcal{B}$). The parties are non-communicating, and therefore cannot coordinate their outputs. The behaviour of the provers is characterized by a correlation, a set of conditional probabilities $\mathsf{p}=\{p(a,b|x,y):a\in\mathcal{A},b\in\mathcal{B},x\in\mathcal{X},y\in\mathcal{Y}\}$, which is realized by an underlying physical theory or model. In the quantum setting, we allow the provers to share a bipartite quantum state, and say the correlation $\mathsf{p}$ is realized by a bipartite (quantum) model
$\mathrm{Q}=\big{(}\mathcal{H}_{A},\mathcal{H}_{B},\{\{M_{a|x}\}_{a\in\mathcal{A}}\}_{x\in\mathcal{X}},\{\{N_{b|y}\}_{b\in\mathcal{B}}\}_{y\in\mathcal{Y}},|\Psi\rangle_{AB}\big{)},$ (2.1)
where $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ are Hilbert spaces, $\{M_{a|x}\}_{a\in\mathcal{A}}$ and $\{N_{b|y}\}_{b\in\mathcal{B}}$ are POVMs on $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ respectively, and $|\Psi\rangle_{AB}$ is a vector state in $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$. More generally, a correlation $\mathsf{p}$ is quantum (or an element of $C_{\mathrm{q}}(n_{A},n_{B},m_{A},m_{B})$) if there exists a bipartite model $\mathrm{Q}$ for which $\mathsf{p}$ can be realized via the Born rule as
$p(a,b,|x,y)=\langle\Psi|M_{a|x}\otimes N_{b|y}|\Psi\rangle.$ (2.2)
We denote the class of bipartite (quantum) models by $\mathcal{Q}(n_{A},n_{B},m_{A},m_{B})$. From now on we will refer to such models simply as bipartite models.
In contrast to the set of quantum correlations, we have the collection of local correlations $C_{\mathrm{loc}}(n_{A},n_{B},m_{A},m_{B})$. These are the correlations $\{p(a,b,|x,y)\}$ for which there exists a classical model, that is a probability distribution $\lambda_{k}$ and a local distributions $p_{k}^{A}(a|x)$ and $p_{k}^{B}(b|y)$ such that
$p(a,b,|x,y)=\sum_{k}\lambda_{k}\,p_{k}^{A}(a|x)\,p_{k}^{B}(b|y).$ (2.3)
We let $\mathrm{C}=(\lambda_{k},\{p_{k}^{A}\},\{p_{k}^{B}\})$ denote a classical model and let $\mathcal{L}(n_{A},n_{B},m_{A},m_{B})$ denote the class of all classical models. In what follows we consider Bell scenarios where $n_{A}=n_{B}=n$, and $m_{A}=m_{B}=m$. Bell’s theorem [Bel64] states that $C_{loc}(2,2)$ is a strict subset of $C_{q}(2,2)$.
Given a Bell scenario $\mathcal{S}$, one can consider a linear (or Bell) functional on the set of correlations
$I=\sum_{a\in\mathcal{A},b\in\mathcal{B},x\in\mathcal{X},y\in\mathcal{Y}}w_{abxy}\,p(a,b,|x,y),$ (2.4)
for coefficients $w_{abxy}\in\mathbb{R}$. A Bell inequality is a functional $I$ and a bound $\eta>0$ such that $I\leq\eta$ for all $\mathsf{p}\in C_{\mathrm{loc}}(n,m)$. Given a functional $I$, the classical value is the maximal value achieved by the classical correlations $\mathsf{p}\in C_{\mathrm{loc}}(n,m)$. We denote this value by $\eta^{\mathrm{L}}\coloneqq\sup_{\mathsf{p}\in C_{\mathrm{loc}}(m,n)}I$. The quantum value for $I$ is the maximal value achieved by the set of quantum correlations $\mathsf{p}\in C_{\mathrm{q}}(m,n)$, and we denote the quantum value on $I$ by $\eta^{\mathrm{Q}}\coloneqq\sup_{\mathsf{p}\in C_{\mathrm{q}}(m,n)}I$. Hence, a Bell violation occurs whenever there is a $\mathsf{p}\in C_{\mathrm{q}}(m,n)$ for which $I>\eta^{\mathrm{L}}$. A violation of a Bell inequality by non-communicating provers employing a quantum model is an indication of entanglement between provers.
Typically if $\eta^{\mathrm{L}}$ is known for a given $I$, the main question is finding an upper bound on $\eta^{\mathrm{Q}}$. In this case, one often considers the Bell operator [^3] For a more mathematically rigorous treatment of such Bell operators and the SOS approach consult [CMM+24]. $S=\sum_{abxy}w_{abxy}\,M_{a|x}N_{b|y}$, and $\langle S\rangle=\langle\Psi|S|\Psi\rangle$ its quantum expectation with respect to $|\Psi\rangle\in\mathcal{H}_{A}\otimes\mathcal{H}_{B}$. Remark that since bipartite models with separable quantum states generate the classical correlations $C_{\mathrm{q}}(m,n)$, $\langle\Psi|S|\Psi\rangle\leq\eta^{\mathrm{L}}$ whenever $|\Psi\rangle$ is separable (unentangled). However, it’s possible that there could be entangled states for which $\langle\Psi^{\prime}|S|\Psi^{\prime}\rangle>\eta^{\mathrm{L}}$. Hence, given a Bell operator $S$, we can recover the maximum classical and quantum values $\eta^{\mathrm{L}}=\sup_{\mathrm{C}\in\mathcal{Q}(n,m)}\langle S\rangle$ and $\eta^{\mathrm{Q}}=\sup_{\mathrm{Q}\in\mathcal{Q}_{n,m}}\langle S\rangle$ respectively. Technically, we have not fixed the dimensions of the Bell operator as we want to consider any finite-dimensional model. Hence, the $\sup$ is implicitly also optimizing over all finite-dimensional Hilbert spaces $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$.
An approach to establishing upper bounds on $\langle S\rangle$ is using sum-of-squares techniques. Let $S$ be a Bell operator and $\eta^{\prime}>0$. The shifted Bell operator $\eta^{\prime}\mathbbm{1}-S$ admits a sum-of-squares (SOS) decomposition if there exists a set of complex polynomials $\{P_{i}\}_{i\in\mathcal{I}}$ in the elements $\{M_{a|x},N_{b|y}:a\in\mathcal{A},b\in\mathcal{B},x\in\mathcal{X},y\in\mathcal{Y}\}$ satisfying
$\eta^{\prime}\mathbbm{1}-S=\sum_{i\in\mathcal{I}}P_{i}^{\dagger}P_{i}.$ (2.5)
The existence of an SOS decomposition for the operator $\eta^{\prime}\mathbbm{1}-S$ implies that $\eta^{\prime}\mathbbm{1}-S$ is positive, and therefore $\eta^{\prime}$ is an upper bound on the maximum quantum value of $\langle S\rangle$. Additionally, if $\eta^{\prime}$ is achievable by a bipartite model, then we write $\eta^{\prime}=\eta^{\mathrm{Q}}$. In this case, the shifted Bell operator is $\bar{S}=\eta^{\mathrm{Q}}\mathbbm{1}-S$, and observing $\langle\Psi|\bar{S}|\Psi\rangle=0$ implies the constraints $P_{i}|\Psi\rangle=0$ for all $i\in\mathcal{I}$; these constraints can often be used to infer the algebraic structure (rigidity) of the measurements $\{M_{a|x}\},\{N_{b|y}\}$ which achieve $\langle S\rangle=\eta^{\mathrm{Q}}$.
3 Compiled Bell inequalities
The compilation procedure of a Bell scenario is essentially the same as the procedure for compiling nonlocal games outlined in [KLVY23]. Let $\mathcal{S}=(\mathcal{X},\mathcal{Y},\mathcal{A},\mathcal{B},\pi)$ be a 2-prover Bell scenario and $\mathsf{Q}$ a quantum homomorphic encryption scheme with security parameter $\lambda$. A compiled Bell scenario is the following 2-round single-prover scenario. To setup, the verifier samples a pair of inputs $(x,y)\in\mathcal{X}\times\mathcal{Y}$ according to the distribution $\pi:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$. Additionally, the verifier samples a secret key $\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})$ and encrypts the first input as the ciphertext $\chi\leftarrow\mathsf{Enc}(\mathsf{sk},x)$.
  1. 1.
    The verifier sends the ciphertext $\chi$ to the prover. The prover replies with a ciphertext $\alpha$ encoding their output. The verifier decrypts obtaining outcome $a\leftarrow\mathsf{Dec}(\mathsf{sk},\alpha)$ from $\mathcal{A}$.
  2. 2.
    The verifier sends the sampled (plaintext) input $y\in\mathcal{Y}$ to the prover, who replies with another outcome $b\in\mathcal{B}$.
In the compiled scenario, the prover acts as a single Hilbert space $\widetilde{\mathcal{H}}$ and prepares an initial state $|\Psi\rangle\in\widetilde{\mathcal{H}}$. Given a secret key $\mathsf{sk}$, the first round of the protocol is characterized by a family of POVMs $\{\{\widetilde{M}_{\alpha|\chi}\}_{\alpha\in\bar{\mathcal{A}}}\}_{\chi\in\bar{\mathcal{X}}}$ and unitaries $\{U_{\alpha,\chi}\}_{\alpha\in\bar{\mathcal{A}},\chi\in\bar{\mathcal{X}}}$, where $\bar{\mathcal{X}}$ and $\bar{\mathcal{A}}$ are the set of all ciphertexts of the first round input and output, respectively. Unlike in the bipartite setting, we must account for unitary operations applied to the post-measurement state in the first round. With this in mind, we denote the sub-normalized post-measurement state given the measurement over ciphertext $\chi$ and encrypted outcome $\alpha$ by
$U_{\alpha,\chi}\widetilde{M}_{\alpha|\chi}|\Psi\rangle\coloneqq|\Psi_{\alpha|\chi}\rangle.$ (3.1)
Note that these vectors are sub-normalized. In particular, the probability of obtaining $\alpha\in\bar{\mathcal{A}}$ given $\chi\in\bar{\mathcal{X}}$ is given by $\langle\Psi_{\alpha|\chi}|\Psi_{\alpha|\chi}\rangle$. In the second round, the device makes a POVM measurement $\{\{N_{b|y}\}_{b\in\mathcal{B}}\}_{y\in\mathcal{Y}}$, where the resulting conditional probability is given by
$\langle\Psi|\widetilde{M}_{\alpha|\chi}^{\dagger}U_{\alpha,\chi}^{\dagger}N_{b|y}U_{\alpha,\chi}\widetilde{M}_{\alpha|\chi}|\Psi\rangle=\langle\Psi_{\alpha|\chi}|N_{b|y}|\Psi_{\alpha|\chi}\rangle,$ (3.2)
for a fixed $\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})$, ciphertexts $\chi\in\bar{\mathcal{X}}$, $\alpha\in\bar{\mathcal{A}}$, and plaintexts $y\in\mathcal{Y}$, $b\in\mathcal{B}$.
To summarize, for a fixed QHE scheme, a compiled (quantum) model is given by a tuple
$\widetilde{\mathrm{Q}}=(\widetilde{\mathcal{H}},\{|\Psi_{\alpha|\chi}\rangle\},\{\{N_{b|y}\}),$ (3.3)
where all the relevant measurements and states are obtained by a QPT procedure. We remark that one can consider a description of the model which includes the initial state $|\Psi\rangle$ and the operators $\{U_{\alpha,\chi}\widetilde{M}_{\alpha|\chi}\}$, rather than the post-measurement states $|\Psi_{\alpha|\chi}\rangle$. Hence, $\widetilde{\mathrm{Q}}$ is really a coarser description of a quantum model in the compiled setting. The joint distribution of the outcomes after both rounds is given by
$p(a,b,|x,y)=\mathop{\mathbb{E}}_{\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})}\mathop{\mathbb{E}}_{\chi:\mathsf{Dec}(\chi)=x}\sum_{\alpha:\mathsf{Dec}(\alpha)=a}\langle\Psi_{\alpha|\chi}|N_{b|y}|\Psi_{\alpha|\chi}\rangle.$ (3.4)
Throughout the work, we will often suppress the expectation over secret keys $\mathop{\mathbb{E}}_{\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})}$ and simply write the expectation for a fixed key.
The marginal distribution $p(a|x)$ is independent of the second input $y$ due to the sequential nature of the protocol. Later, we will need to establish a type of independence between the second outcome $b$ and the first input $x$. To do so we will need to consider the distributions of the decrypted outputs and appeal to the security promise of the QHE scheme. Specifically, we require a key lemma which has appeared in several works [NZ23, CMM+24]. We borrow a version from [BVB+24], which we refer the reader to for proof.
Lemma 3.1 ([BVB+24], Lemma 7).
Let $\lambda\in\mathbb{N}$ be the security parameter for a QHE scheme. There exists a negligible function $\delta_{\mathsf{QHE}}$ of $\lambda$ such that the following holds. For any two QPT sampleable distributions $\mathcal{D}_{1},\mathcal{D}_{2}$ over plaintext inputs $\mathcal{A}$, QPT preparable state $|\Psi\rangle$, and QPT implementable operation $\mathcal{N}$, it holds that
$\left|\mathop{\mathbb{E}}_{x\leftarrow\mathcal{D}_{1}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\mathsf{Enc}(x)=\chi\end{subarray}}\sum_{\alpha,\beta}\langle\Psi_{\alpha|\chi}|\mathcal{N}|\Psi_{\alpha|\chi}\rangle-\mathop{\mathbb{E}}_{x\leftarrow\mathcal{D}_{2}}\mathop{\mathbb{E}}_{\begin{subarray}{c}\mathsf{Enc}(x)=\chi\end{subarray}}\sum_{\alpha,\beta}\langle\Psi_{\alpha|\chi}|\mathcal{N}|\Psi_{\alpha|\chi}\rangle\right|\leq\delta_{\mathsf{QHE}}.$
The approximate no-signalling conditions from Alice to Bob can then be seen by applying Lemma 3.1, since
$\Big{|}\mathop{\mathbb{E}}_{\chi:\mathsf{Dec}(\chi)=x}\sum_{\alpha\in\bar{\mathcal{A}}}\langle\Psi_{\alpha|\chi}|N_{b|y}|\Psi_{\alpha|\chi}\rangle-\mathop{\mathbb{E}}_{\chi:\mathsf{Dec}(\chi)=x^{\prime}}\sum_{\alpha\in\bar{\mathcal{A}}}\langle\Psi_{\alpha|\chi}|N_{b|y}|\Psi_{\alpha|\chi}\rangle\Big{|}\leq\delta_{\mathsf{QHE}}$ (3.5)
holds for all $b\in\mathcal{B},y\in\mathcal{Y}$ and $x,x^{\prime}\in\mathcal{X}$ with $x\neq x^{\prime}$.
3.1 Quantum bounds for compiled inequalities
A compiled (quantum) model describes the correlation $\{p(a,b,|x,y)\}$ observed in a compiled Bell scenario. A compiled Bell functional is a linear functional $I$ evaluated on correlations realized by compiled models. That is
$I=\sum_{abxy}w_{abxy}\mathop{\mathbb{E}}_{\begin{subarray}{c}\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})\\ \chi:\mathsf{Dec}(\chi)=x\end{subarray}}\sum_{\alpha:\mathsf{Dec}(\alpha)=a}\langle\Psi_{\alpha|\chi}|N_{b|y}|\Psi_{\alpha|\chi}\rangle.$ (3.6)
The compilation procedure preserves the classical bound of the Bell inequality (up to negligible error)[KLVY23]. Hence in the compiled setting, classical provers cannot effectively violate a Bell inequality by more than they could in the classical (bipartite) scenario.
We now turn our attention to quantum provers in the compiled setting. The results of [KLVY23] imply that a quantum prover in the compiled setting can achieve the maximal bound in the bipartite setting. However, the existence of a quantum compiled behaviour which exceeds the maximal quantum bell violation in the bipartite case (by more than negligible factors) has not been ruled out. Nonetheless, in several cases (like the CHSH inequality and more generally all XOR games [CMM+24]) we know that the quantum compiled behaviour cannot exceed the value $\eta^{\mathrm{Q}}$ by more than negligible amounts. One technique for establishing such bounds was introduced in [NZ23] and uses SOS techniques to bound the quantum violation of the compiled Bell functional.
3.2 Extended definition of the pseudo-expectation
Our approach builds off the methods used in [NZ23] and [CMM+24] using pseudo-expectations. This SOS approach necessitates the definition of a pseudo-expectation; defined as a unital, linear map from a subspace $\mathcal{T}$ of the algebra generated by $\{M_{a|x},N_{b|y}\}_{a\in\mathcal{A},x\in\mathcal{X},b\in\mathcal{B},y\in\mathcal{Y}}$ to the complex numbers $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}:\mathcal{T}\mapsto\mathbb{C}$. In the case $n=m=2$, it suffices to define the pseudo expectation $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}$ on the observables $A_{x}=\sum_{a\in\{0,1\}}(-1)^{a}M_{a|x},\ B_{y}=\sum_{b\in\{0,1\}}(-1)^{b}N_{b|y}$ and require that they are mapped to their expectations in the compiled scenario [^4] Though in the following we define $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}$ for $n=m=2$, this can be directly extended to arbitrary Bell scenarios by defining $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}$ on the POVM elements $M_{a|x},N_{b|y}$ in an analogous way.. We also require $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}$ to be approximately positive on squares; that is, for all $w\in\mathcal{T}$, $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[w^{\dagger}w]\geq-\delta$ for some $\delta\geq 0$. In previous works, the definition of the pseudo-expectation has been restricted to at most degree 2 monomials, consisting of at most one Alice and one Bob observable:
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[A_{x}B_{y}]$ $:=\mathop{\mathbb{E}}_{\chi:\mathsf{Dec}(\chi)=x}\sum_{\alpha}(-1)^{\mathsf{Dec}(\alpha)}\langle\Psi_{\alpha|\chi}|B_{y}|\Psi_{\alpha|\chi}\rangle,$ (3.7)
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[A_{x}A_{x^{\prime}}]$ $:=\delta_{x,x^{\prime}},$
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[B_{y}B_{y^{\prime}}]$ $:=\mathop{\mathbb{E}}_{x\in\mathcal{X}}\mathop{\mathbb{E}}_{\chi:\mathsf{Dec}(\chi)=x}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|B_{y}B_{y^{\prime}}|\Psi_{\alpha|\chi}\rangle,$
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[A_{x}]$ $:=\mathop{\mathbb{E}}_{\chi:\mathsf{Dec}(\chi)=x}\sum_{\alpha}(-1)^{\mathsf{Dec}(\alpha)}\langle\Psi_{\alpha|\chi}|\Psi_{\alpha|\chi}\rangle$
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[B_{y}]$ $:=\mathop{\mathbb{E}}_{x\in\mathcal{X}}\mathop{\mathbb{E}}_{\chi:\mathsf{Dec}(\chi)=x}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|B_{y}|\Psi_{\alpha|\chi}\rangle,$
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\mathbbm{1}]$ $:=1,$
where $\mathop{\mathbb{E}}_{x\in\mathcal{X}}$ denotes the expectation according to an arbitrary fixed distribution over $\mathcal{X}$. This is already sufficient to handle known SOS decompositions for a variety of well-studied Bell inequalities whenever the polynomials are expressed in the basis $\{\mathbbm{1},A_{x},B_{y}\}_{x,y}$. However, there are Bell inequalities, such as the tilted-CHSH inequality, for which no known SOS decomposition exists in the basis $\{\mathbbm{1},A_{x},B_{y}\}_{x,y}$.
The contribution of this section is to expand the definition of the pseudo-expectation to the basis encompassing all monomials in $A_{x},B_{0},B_{1}$, for a fixed $x\in\mathcal{X}$. This allows us to handle more general SOS decompositions, and in particular, the tilted-CHSH inequalities. Let $w(A_{x},B_{0},B_{1})$ be a monomial in the elements $\{A_{x},B_{0},B_{1}\}$. Importantly, $x$ is fixed, and we do not consider monomials of the form $A_{0}A_{1}B_{y}$. Let $\bar{w}$ be the canonical form of $w$ under the relations $[A_{x},B_{y}]=0$, $(B_{y})^{2}=(A_{x})^{2}=\mathbbm{1}$, where all $A_{x}$ terms are commuted to the left. Since we only consider one value of $x$, these will all be of the form $(A_{x})^{i}\bar{w}(B_{0},B_{1})$ for some $i\in\{0,1\}$, where the monomial $\bar{w}(B_{0},B_{1})$ cannot be reduced further. We then define the pseudo-expectation
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}\big{[}w(A_{x},B_{0},B_{1})\big{]}:=\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}\big{[}(A_{x})^{i}\bar{w}(B_{0},B_{1})\big{]}.$ (3.8)
For the case $i=0$, we define
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}\big{[}\bar{w}(B_{0},B_{1})\big{]}:=\mathop{\mathbb{E}}_{x\in\mathcal{X}}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\bar{w}(B_{0},B_{1})|\Psi_{\alpha|\chi}\rangle,$ (3.9)
and for the case $i=1$,
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}\big{[}A_{x}\bar{w}(B_{0},B_{1})\big{]}:=\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}(-1)^{\mathsf{Dec}(\alpha)}\langle\Psi_{\alpha|\chi}|\bar{w}(B_{0},B_{1})|\Psi_{\alpha|\chi}\rangle.$ (3.10)
From the above definitions, we state the main result of this section in the following lemma, which can be applied generally to any polynomial expressible in the basis $\{A_{x},B_{0},B_{1}\}$.
Lemma 3.2.
Let $A_{x}$ and $\{B_{y}\}_{y\in\mathcal{Y}}$ be binary observables, and let
$P=\sum_{i}\gamma_{i}(A_{x})^{k_{i}}w_{i}(B_{0},B_{1}),$ (3.11)
where $\gamma_{i}\in\mathbb{C}$, $k_{i}\in\{0,1\}$ and $w_{i}(B_{0},B_{1})$ is any monomial in the algebra of $\{B_{0},B_{1}\}$. Then there exists a negligible function $\delta_{\mathsf{QHE}}$ of the security parameter $\lambda\in\mathbb{N}$ such that
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[P^{\dagger}P]\geq-\delta_{\mathsf{QHE}}.$ (3.12)
Proof.
To begin, we write
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[P^{\dagger}P]$ $=\sum_{ij}\gamma_{i}^{*}\gamma_{j}\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[(A_{x})^{k_{i}}w_{i}(B_{0},B_{1})(A_{x})^{k_{j}}w_{j}(B_{0},B_{1})]$ (3.13)
$=\sum_{ij}\gamma_{i}^{*}\gamma_{j}\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[(A_{x})^{k_{i}+k_{j}}\bar{w}_{ij}(B_{0},B_{1})],$
where we used the linearity of $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\cdot]$ in the first line, and in the second line we used the fact that $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[w]=\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\bar{w}]$ (where $\bar{w}$ is the canonical form of the monomial $w$), and defined $\bar{w}_{ij}$ to be the canonical form of $w_{i}w_{j}$. We now need to consider two types of terms. First, when $k_{i}\oplus k_{j}=0$, we apply the definition in Eq. 3.9 in conjunction with Lemma 3.1 to write
$\sum_{ij:k_{i}\oplus k_{j}=0}\gamma_{i}^{*}\gamma_{j}\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\bar{w}_{ij}(B_{0},B_{1})]$ (3.14)
$=\mathop{\mathbb{E}}_{x^{\prime}\in\mathcal{X}}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x^{\prime})=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\Bigg{(}\sum_{ij:k_{i}\oplus k_{j}=0}\gamma_{i}^{*}\gamma_{j}\bar{w}_{ij}(B_{0},B_{1})\Bigg{)}|\Psi_{\alpha|\chi}\rangle$
$\approx_{\delta_{\mathsf{QHE}}}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\Bigg{(}\sum_{ij:k_{i}\oplus k_{j}=0}\gamma_{i}^{*}\gamma_{j}\bar{w}_{ij}(B_{0},B_{1})\Bigg{)}|\Psi_{\alpha|\chi}\rangle$
$=\sum_{ij:k_{i}\oplus k_{j}=0}\gamma_{i}^{*}\gamma_{j}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\bar{w}_{ij}(B_{0},B_{1})|\Psi_{\alpha|\chi}\rangle.$
When $k_{i}\oplus k_{j}=1$, we can apply Eq. 3.10 directly. Putting these two together, we observe
$\sum_{ij}\gamma_{i}^{*}\gamma_{j}\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[(A_{x})^{k_{i}+k_{j}}\bar{w}_{ij}(B_{0},B_{1})]$ (3.15)
$=\sum_{ij:k_{i}\oplus k_{j}=0}\gamma_{i}^{*}\gamma_{j}\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\bar{w}_{ij}]+\sum_{ij:k_{i}\oplus k_{j}=1}\gamma_{i}^{*}\gamma_{j}\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[A_{x}\bar{w}_{ij}]$
$\approx_{\delta_{\mathsf{QHE}}}\sum_{ij:k_{i}\oplus k_{j}=0}\gamma_{i}^{*}\gamma_{j}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\bar{w}_{ij}(B_{0},B_{1})|\Psi_{\alpha|\chi}\rangle$
$+\sum_{ij:k_{i}\oplus k_{j}=1}\gamma_{i}^{*}\gamma_{j}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}(-1)^{\mathsf{Dec}(\alpha)}\langle\Psi_{\alpha|\chi}|\bar{w}_{ij}(B_{0},B_{1})|\Psi_{\alpha|\chi}\rangle$
$=\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\sum_{ij}(-1)^{\mathsf{Dec}(\alpha)\cdot(k_{i}+k_{j})}\gamma_{i}^{*}\gamma_{j}w_{i}(B_{0},B_{1})w_{j}(B_{0},B_{1})|\Psi_{\alpha|\chi}\rangle$
$=\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\Big{|}\sum_{i}(-1)^{\mathsf{Dec}(\alpha)\cdot k_{i}}\gamma_{i}w_{i}(B_{0},B_{1})\Big{|}^{2}|\Psi_{\alpha|\chi}\rangle\geq 0.$
Where in the fifth line, we used the fact that Bob’s observables satisfy the canonical relations, so we can always replace the canonical monomial $\bar{w}_{ij}$ with $w_{i}w_{j}$. The final line is obtained by noting the square inside the expectation. We therefore conclude $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[P^{\dagger}P]\approx_{\delta_{\mathsf{QHE}}}h$ for some $h\geq 0$, which implies $|\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[P^{\dagger}P]-h|\leq\delta_{\mathsf{QHE}}$, and $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[P^{\dagger}P]\geq h-\delta_{\mathsf{QHE}}\geq-\delta_{\mathsf{QHE}}$ as required. ∎
3.3 Self-testing in the compiled setting
In the compiled setting, a desirable self-testing result would imply that any compiled model achieving the optimal correlation would be efficiently isometric to a canonical compiled model. Furthermore, in cases where the quantum compiled value equals the bipartite value, one would expect isometries relating each compiled model achieving the quantum value to the ideal (bipartite) quantum model.
One of the main obstructions to obtaining strong self-testing results in the compiled setting is the lack of rigidity for the quantum operations applied under the QHE. Nevertheless, in the following we show there are mathematical techniques used to establish self-testing results in the bipartite setting which still provide some partial self-testing results in the compiled setting.
To begin, we recall the description of a bipartite model
$\mathrm{Q}=\big{(}\mathcal{H}_{A},\mathcal{H}_{B},\{\{P_{a|x}\}_{a\in\mathcal{A}}\}_{x\in\mathcal{X}},\{\{Q_{b|y}\}_{b\in\mathcal{B}}\}_{y\in\mathcal{Y}},|\phi\rangle\big{)},$ (3.16)
and consider the set of (sub-normalized) post-measurement states obtained after Alice has measured and $\mathcal{H}_{A}$ is traced out. That is,
$\{\rho_{a|x}:a\in\mathcal{A},x\in\mathcal{X}\}:=\{{\mathrm{t}r}_{A}[(P_{a|x}\otimes\mathbbm{1})|\phi\rangle\langle\phi|]:a\in\mathcal{A},x\in\mathcal{X}\}.$ (3.17)
Definition 3.3.
A bipartite model $\mathrm{Q}$ is said to have a pure $B$-model if $\rho_{a|x}$ is a pure for all $a\in\mathcal{A}$, $x\in\mathcal{X}$. In this case, we write the reduced states as pure states $|\varphi_{a|x}\rangle\in\mathcal{H}_{B}$ satisfying $\rho_{a|x}=|\varphi_{a|x}\rangle\!\langle\varphi_{a|x}|$.
Based on the above, we denote the compiled counter-part of a pure $B$-model $\mathrm{Q}$ as
$\widetilde{\mathrm{Q}}^{*}|_{B}=\big{(}\mathcal{H}_{B},\{|\widetilde{\phi}_{\alpha|\chi}\rangle\}_{\alpha\in\bar{\mathcal{A}},\chi\in\bar{\mathcal{X}}},\{\{Q_{b|y}\}_{b\in\mathcal{B}}\}_{y\in\mathcal{Y}}\big{)},$ (3.18)
where the sub-normalized states $|\widetilde{\phi}_{\alpha|\chi}\rangle$ are given by the relation
$|\widetilde{\phi}_{\alpha|\chi}\rangle:=|\varphi_{a|x}\rangle,\ \forall\chi:\mathsf{Enc}(x)=\chi,\ \forall\alpha:\mathsf{Dec}(\alpha)=a.$ (3.19)
Note that the above implies
$p(a,b,|x,y)$ $=\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha:\mathsf{Dec}(\alpha)=a}\langle\widetilde{\phi}_{\alpha|\chi}|Q_{b|y}|\widetilde{\phi}_{\alpha|\chi}\rangle$ (3.20)
$={\mathrm{t}r}\big{[}(P_{a|x}\otimes Q_{b|y})|\phi\rangle\!\langle\phi|\big{]},$
where we used the inequality ${\mathrm{t}r}_{A}[X_{AB}]Y_{B}={\mathrm{t}r}_{A}[X_{AB}(\mathbbm{1}\otimes Y_{B})]$, telling us that the model $\widetilde{\mathrm{Q}}^{*}|_{B}$ satisfies $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S]=\eta^{\mathrm{Q}}$.
Based on these definitions, we present a notion of self-testing for compiled correlations. Consider a Bell functional $I$ in the bipartite scenario with Bell operator $S$, a maximum bipartite value $\eta^{\mathrm{Q}}$, and ideal bipartite model $\mathrm{Q}^{*}$. Further suppose that $\mathrm{Q}^{*}$ has compiled counter-part $\widetilde{\mathrm{Q}}^{*}|_{B}$.
Definition 3.4.
The inequality $I\leq\eta^{\mathrm{Q}}$ is a compiled self-test for $\widetilde{\mathrm{Q}}^{*}|_{B}$ if, for every compiled model $\widetilde{\mathrm{Q}}$ which achieves $I\geq\eta^{\mathrm{Q}}-\epsilon$, there exists a negligible function $\delta_{\mathsf{QHE}}$, a function $f(\epsilon)$ such that $f\rightarrow 0$ as $\epsilon\rightarrow 0$, an efficient isometry $V:\widetilde{\mathcal{H}}\mapsto\mathbb{C}^{n}\otimes\mathcal{H}_{\mathrm{aux}}$, and auxiliary states $\{|\mathsf{aux}_{\alpha|\chi}\rangle\}_{\alpha\in\bar{\mathcal{A}},\chi\in\bar{\mathcal{X}}}\subset\mathcal{H}_{\mathrm{aux}}$, which satisfy
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\big{\|}VN_{b|y}|\Psi_{\alpha|\chi}\rangle-Q_{b|y}|\varphi_{\mathsf{Dec}(\alpha)|x}\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big{\|}^{2}\leq f(\epsilon)+\delta_{\mathsf{QHE}},$ (3.21)
for all $x\in\mathcal{X},b\in\mathcal{B}$, and $y\in\mathcal{Y}$.
Notice that satisfying Eq. 3.21 implies self-testing of the post-measurement states $\{|\Psi_{\alpha|\chi}\rangle\}$ alone. This can be seen by summing both sides over $b\in\mathcal{B}$ and applying the triangle inequality. Although we do not expect to achieve zero error in the compiled setting, we can in theory consider the case where $\epsilon=\delta_{\mathsf{QHE}}=0$. In this case Eq. 3.21 would imply that
$VN_{b|y}|\Psi_{\alpha|\chi}\rangle\!\langle\Psi_{\alpha|\chi}|N_{b|y}V^{\dagger}=Q_{b|y}{\mathrm{t}r}_{A}\big{[}(P_{a|x}\otimes\mathbbm{1}_{B})|\phi\rangle\!\langle\phi|]Q_{b|y}\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\!\langle\mathsf{aux}_{\alpha|\chi}|,$ (3.22)
for all $\alpha$ and $\chi$ such that $a=\mathsf{Dec}(\alpha)$, $\chi=\mathsf{Enc}(x)$. In other words, for all compiled models which maximally violate the Bell expression, we observe that the state after the first round is isometric to preparing a pure bipartite entangled state $|\phi\rangle$ and measuring the first subsystem, according to the honest (ideal) model.
Remark 3.5.
Suppose the ideal nonlocal model employs the maximally entangled state $(|00\rangle+|11\rangle)/\sqrt{2}$, and Alice’s measurements include $A_{0}=\sigma_{Z}$ (as is the case for e.g. the CHSH inequality). Then, in the compiled setting, the honest post-measurement state on Bob’s register after Alice measures $A_{0}$ and gets output $a$ is $|\varphi_{\mathsf{Dec}(\alpha)|0}\rangle=|\mathsf{Dec}(\alpha)\rangle/\sqrt{2}$. Now consider a dishonest device that does not maximally violate the Bell inequality; can we still satisfy Eq. 3.21? One could consider a trivial isometry that ignores the state $|\Psi_{\alpha|\chi}\rangle$ and deterministically prepares $|0\rangle$, via $V|\Psi_{\alpha|\chi}\rangle=(1/\sqrt{2})|0\rangle\otimes|\hat{\Psi}_{\alpha|\chi}\rangle$, where $|\hat{\Psi}_{\alpha|\chi}\rangle=\sqrt{2}|\Psi_{\alpha|\chi}\rangle$ is normalized. Then
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(0)=\chi}\sum_{\alpha}\|V|\Psi_{\alpha|\chi}\rangle-|\varphi_{\mathsf{Dec}(\alpha)|0}\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\|^{2}$ (3.23)
$=\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(0)=\chi}\sum_{\alpha}\frac{1}{2}\||0\rangle-|\mathsf{Dec}(\alpha)\rangle\|^{2}=1.$
Intuitively, any isometry preparing a fixed state independent of $\alpha$ will result in a large distance after summing over the ciphertexts $\alpha$, if the states $|\varphi_{a|x}\rangle$ are orthogonal. This forces the isometry to prepare a state dependent on $\alpha$. Moreover, it must do this according to $\mathsf{Dec}(\alpha)$ to be successful, which would seem to violate the encryption.
4 Tilted-CHSH is a compiled self-test
In this section, we show that the extended tilted-CHSH inequalities satisfy our notion of a compiled self-test.
4.1 Quantum bounds for tilted-CHSH expressions
We now present the family of tilted-CHSH type expressions and their SOS decompositions discovered in [BSB24]. Let $\theta\in(0,\pi/4]$, $\phi\in\big{(}\max\{-2\theta,-\pi+2\theta\},\min\{2\theta,\pi-2\theta\}\big{)}\setminus\{0\}$, and $\lambda_{\theta,\phi}\in\mathbb{R}$ such that
$\frac{1}{\lambda_{\theta,\phi}^{2}}=\frac{\sin^{2}(2\theta)}{\tan^{2}(\phi)}-\cos^{2}(2\theta).$ (4.1)
From here, we define the following expressions:
$S_{\theta,\phi}$ $:=A_{0}\frac{B_{0}+B_{1}}{\cos(\phi)}+\lambda_{\theta,\phi}^{2}\Big{[}\sin(2\theta)A_{1}\frac{B_{0}-B_{1}}{\sin(\phi)}+\cos(2\theta)\frac{B_{0}+B_{1}}{\cos(\phi)}\Big{]},$ (4.2)
$\eta^{\mathrm{Q}}_{\theta,\phi}$ $:=2(1+\lambda_{\theta,\phi}^{2}).$
We also let $I_{\theta,\phi}$ denote the corresponding Bell functional, and recall the following result.
Lemma 4.1 ([BSB24], Section 3.2.1).
Let $\theta\in(0,\pi/4]$, $\phi\in\big{(}\max\{-2\theta,-\pi+2\theta\},\min\{2\theta,\pi-2\theta\}\big{)}\setminus\{0\}$, $\lambda_{\theta,\phi}$ be given by Eq. 4.1 and $S_{\theta,\phi},\eta^{\mathrm{Q}}_{\theta,\phi}$ be defined in Eq. 4.2. Define the following polynomials:
$N_{0}$ $:=A_{0}-\frac{B_{0}+B_{1}}{2\cos(\phi)},$ (4.3)
$N_{1}$ $:=A_{1}-\sin(2\theta)\frac{B_{0}-B_{1}}{2\sin(\phi)}-\cos(2\theta)A_{1}\frac{B_{0}+B_{1}}{2\cos(\phi)}.$
The shifted Bell operator $\bar{S}_{\theta,\phi}=\eta^{\mathrm{Q}}_{\theta,\phi}\mathbbm{1}-S_{\theta,\phi}$ admits the SOS decomposition
$\bar{S}_{\theta,\phi}=N_{0}^{\dagger}N_{0}+\lambda_{\theta,\phi}^{2}N_{1}^{\dagger}N_{1}.$ (4.4)
Using the decomposition in Eq. 4.4, it was shown in [BSB24] that the inequality $\langle S_{\theta,\phi}\rangle\leq\eta_{\theta,\phi}^{\mathrm{Q}}$ self-tests the partially entangled state $|\psi_{\theta}\rangle=\cos(\theta)|00\rangle+\sin(\theta)|11\rangle$ and the measurements
$A_{0}$ $=\sigma_{Z},\ A_{1}=\sigma_{X},$ (4.5)
$B_{y}$ $=\cos(\phi)\,\sigma_{Z}+(-1)^{y}\sin(\phi)\,\sigma_{X},\ y\in\{0,1\},$
where $\sigma_{Z},\sigma_{X}$ are the Pauli operators. Notably, by setting $\phi=\mu_{\theta}$ where $\tan(\mu_{\theta})=\sin(2\theta)$, this family encompasses what are most commonly referred to as “tilted-CHSH inequalities” given by the Bell operator
$T_{\theta}=\alpha_{\theta}A_{0}+A_{0}(B_{0}+B_{1})+A_{1}(B_{0}-B_{1}),$ (4.6)
where $\alpha_{\theta}=2/\sqrt{1+2\tan^{2}(2\theta)}$  [AMP12, YN13, BP15]. Compared to the SOS decompositions for $T_{\theta}$ from [BP15], the decomposition of [BSB24] is expressed in the basis for which our extended pseudo-expectation is well defined (cf. Eq. 3.12), allowing us to provide bounds on the compiled value of $T_{\theta}$, and more generally the family $S_{\theta,\phi}$.
Theorem 4.2.
Let $\theta\in(0,\pi/4]$, $\phi\in\big{(}\max\{-2\theta,-\pi+2\theta\},\min\{2\theta,\pi-2\theta\}\big{)}\setminus\{0\}$, and let $S_{\theta,\phi}$ be the extended tilted-CHSH expression with quantum bound $\eta^{\mathrm{Q}}_{\theta,\phi}$, given by Eq. 4.2. Then the maximum quantum value of the corresponding compiled Bell inequality is given by $\eta^{\mathrm{Q}}_{\theta,\phi}+\delta_{\mathsf{QHE}}^{\prime}$, where $\delta_{\mathsf{QHE}}^{\prime}$ is a negligible function of the security parameter.
Proof.
We evaluate the pseudo-expectation on the shifted Bell expression $\bar{S}_{\theta,\phi}$:
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\bar{S}_{\theta,\phi}]$ $=\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{0}^{\dagger}N_{0}]+\lambda^{2}_{\theta,\phi}\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{1}^{\dagger}N_{1}],$ (4.7)
where we used the decomposition in Eq. 4.4. The polynomial $N_{0}$ is expressed in the basis $\{A_{0},B_{0},B_{1}\}$, and we find by Eq. 3.12 that
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{0}^{\dagger}N_{0}]\geq-\delta_{\mathsf{QHE}}.$ (4.8)
Similarly, $N_{1}$ is expressed in the basis $\{A_{1},B_{0},B_{1}\}$, and we see by Eq. 3.12 that $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{1}^{\dagger}N_{1}]\geq-\delta_{\mathsf{QHE}}$. Putting these together, we obtain
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\bar{S}_{\theta,\phi}]\geq-\delta_{\mathsf{QHE}}(1+\lambda^{2}_{\theta,\phi})=:-\delta_{\mathsf{QHE}}^{\prime},$ (4.9)
which implies $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S_{\theta,\phi}]\leq\eta^{\mathrm{Q}}_{\theta,\phi}+\delta_{\mathsf{QHE}}^{\prime}$ as desired, where $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S_{\theta,\phi}]$ is the expected value of the compiled Bell inequality. ∎
Remark 4.3.
The extension of the $S_{\theta,\phi}$ family presented in [[BSB24], Section 3.2.3] self-tests the state $|\psi_{\theta}\rangle$ along with the more general measurements
$A_{0}$ $=\sigma_{Z},\ A_{1}=\sigma_{X},$ (4.10)
$B_{0}$ $=\cos(\phi)\,\sigma_{Z}+\sin(\phi)\,\sigma_{X},$
$B_{1}$ $=\cos(\omega)\,\sigma_{Z}+\sin(\omega)\,\sigma_{X},$
for $\phi\in(-2\theta,0)$ and $\omega\in(0,2\theta)$. This family of Bell inequalities can also be compiled under our definition of the pseudo-expectation. This is because each SOS polynomial is given in the basis $\{A_{x},B_{0},B_{1}\}$ for a fixed $x$, and we can apply Eq. 3.12 directly as was done in Theorem 4.2. We omit the explicit proof of this for brevity.
4.2 Compiled self-test for tilted-CHSH inequalities
We now show that the extended tilted-CHSH Bell expressions are a compiled self-test according to Eq. 3.21. To do so, we use the fact that the quantum bound is preserved under the compilation procedure (cf. Theorem 4.2), along with the SOS decomposition given by [BSB24] to derive a bound of the form in Eq. 3.21 via a series of triangle inequalities. Our approach follows closely that of Bamps and Pironio [BP15], who obtained robust self-testing bounds for the Bell expression $T_{\theta}$ in the nonlocal scenario, and includes similar calculations to those used in [NZ23, BVB+24], which establish rigidity statements in the compiled setting.
Theorem 4.4.
Let $\theta\in(0,\pi/4]$, $\phi\in\big{(}\max\{-2\theta,-\pi+2\theta\},\min\{2\theta,\pi-2\theta\}\big{)}\setminus\{0\}$, and let $I_{\theta,\phi}$ be the generalized tilted-CHSH functional with quantum bound $\eta^{\mathrm{Q}}_{\theta,\phi}$, given by Eq. 4.2. Then the inequality $I_{\theta,\phi}\leq\eta^{\mathrm{Q}}_{\theta,\phi}$ is a compiled self-test according to Eq. 3.21.
The remainder of this section is devoted to proving the above theorem. Recall the SOS decomposition for $\bar{S}_{\theta,\phi}$, given by the polynomials $N_{0},N_{1}$ defined in Eq. 4.3. Then for compiled models $I_{\theta,\phi}=\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S_{\theta,\phi}]\geq\eta^{\mathrm{Q}}_{\theta,\phi}-\epsilon$, for some $\epsilon\geq 0$, implies $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\bar{S}_{\theta,\phi}]=\eta^{\mathrm{Q}}_{\theta,\phi}-\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S_{\theta,\phi}]\leq\epsilon$. Using the calculation in Eq. 3.15 with $P=N_{0}$ and $P=N_{1}$ respectively, we find
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{0}^{\dagger}N_{0}]$ $\approx_{\delta_{\mathsf{QHE}}}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\Big{|}(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}-Z_{B}\Big{|}^{2}|\Psi_{\alpha|\chi}\rangle=:\hat{E}_{0},$ (4.11)
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{1}^{\dagger}N_{1}]$ $\approx_{\delta_{\mathsf{QHE}}}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\Big{|}\mathbbm{1}-(-1)^{\mathsf{Dec}(\alpha)}\sin(2\theta)X_{B}-\cos(2\theta)Z_{B}\Big{|}^{2}|\Psi_{\alpha|\chi}\rangle=:\hat{E}_{1},$
where we defined $Z_{B}:=(B_{0}+B_{1})/(2\cos(\phi))$ and $X_{B}:=(B_{0}-B_{1})/(2\sin(\phi))$, and note that $\hat{E}_{i}\geq 0$ for $i=0,1$. This implies $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{i}^{\dagger}N_{i}]\geq\hat{E}_{i}-\delta_{\mathsf{QHE}}$, from which we observe that
$\hat{E}_{0}+\lambda^{2}_{\theta,\phi}\hat{E}_{1}\leq\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{0}^{\dagger}N_{0}]+\lambda^{2}_{\theta,\phi}\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{1}^{\dagger}N_{1}]+\delta^{\prime}_{\mathsf{QHE}}=\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\bar{S}_{\theta,\phi}]+\delta^{\prime}_{\mathsf{QHE}}\leq\epsilon+\delta^{\prime}_{\mathsf{QHE}},$ (4.12)
where $\delta^{\prime}_{\mathsf{QHE}}=(1+\lambda^{2}_{\theta,\phi})\delta_{\mathsf{QHE}}$. Since $\hat{E}_{i}\geq 0$, we arrive at
$\hat{E}_{0}\leq\epsilon+\delta^{\prime}_{\mathsf{QHE}}\ \ \mathrm{and}\ \ \lambda^{2}_{\theta,\phi}\hat{E}_{1}\leq\epsilon+\delta^{\prime}_{\mathsf{QHE}}.$ (4.13)
The next part of the proof proceeds as follows: using each constraint in Eq. 4.13, we deduce the (approximate) algebraic structure of the operators $Z_{B}$ and $X_{B}$ on the states $|\Psi_{\alpha|\chi}\rangle$. Namely, we establish anti-commutation relations and approximate unitary relations. We break the proof up into two sections: extracting the states, and extracting the measurements.
Extracting the states: We begin by studying the relations imposed by $\hat{E}_{0}\leq\epsilon+\delta^{\prime}_{\mathsf{QHE}}$. Define $\delta_{0}:=\epsilon+\delta^{\prime}_{\mathsf{QHE}}$. We begin with several claims:
Claim 1.
$\hat{E}_{0}=\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\big{\|}\big{(}(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}-Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{0}$.
Claim 2.
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\big{\|}\big{(}\mathbbm{1}-(Z_{B})^{2}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{0}\big{(}1+1/(2\cos^{2}(\phi))\big{)}=:\delta_{1}$.
Proof.
Note that
$\big{\|}\big{(}\mathbbm{1}-(Z_{B})^{2}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$ $\leq\big{\|}Z_{B}\big{(}(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}-Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\ \ \ \ +\big{\|}(-1)^{\mathsf{Dec}(\alpha)}\big{(}(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}-Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\leq(1+\|Z_{B}\|_{\mathrm{op}}^{2})\big{\|}\big{(}(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}-Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
In the first inequality, we used the triangle inequality and the bound on $\hat{E}_{0}$, and for the second inequality, we used the definition of the operator norm $\|A|\psi\rangle\|^{2}\leq\|A\|_{\mathrm{op}}^{2}\||\psi\rangle\|^{2},\ \ \forall\;|\psi\rangle\in\mathcal{H}$. By the definition of $Z_{B}$, $\|Z_{B}\|_{\mathrm{op}}^{2}\leq(1/(4\cos^{2}(\phi)))(\|B_{0}\|_{\mathrm{op}}^{2}+\|B_{1}\|_{\mathrm{op}}^{2})=1/(2\cos^{2}(\phi))$. Taking the expectation over $\chi:\mathsf{Enc}(x=0)=\chi$ and summing over $\alpha$, we apply 1 to get the desired bound. ∎
Claim 3.
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\big{\|}\big{(}2\cos(2\phi)\mathbbm{1}-\{B_{0},B_{1}\}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq 16\cos^{4}(\phi)\delta_{1}=:\delta_{2}$.
Proof.
Note that $(Z_{B})^{2}=(2\mathbbm{1}+\{B_{0},B_{1}\})/(4\cos^{2}(\phi))$, and therefore
$\big{\|}\big{(}\mathbbm{1}-(Z_{B})^{2}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$ $=\big{\|}\big{(}\mathbbm{1}-(2\mathbbm{1}+\{B_{0},B_{1}\})/(4\cos^{2}(\phi))\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$=\frac{1}{16\cos^{4}(\phi)}\big{\|}\big{(}2\cos(2\phi)\mathbbm{1}-\{B_{0},B_{1}\}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
Taking the expectation over $\chi:\mathsf{Enc}(x=0)=\chi$ and summing over $\alpha$, followed by applying 2, establishes the desired bound. ∎
Claim 4.
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\big{\|}\big{(}\mathbbm{1}-(X_{B})^{2}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{1}/\tan^{4}(\phi)=:\delta_{3}$.
Proof.
Noting $(X_{B})^{2}=(2\mathbbm{1}-\{B_{0},B_{1}\})/(4\sin^{2}(\phi))$,
$\big{\|}\big{(}\mathbbm{1}-(X_{B})^{2}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$ $=\big{\|}\big{(}\mathbbm{1}-(2\mathbbm{1}-\{B_{0},B_{1}\})/(4\sin^{2}(\phi))\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$=\frac{1}{16\sin^{4}(\phi)}\big{\|}\big{(}2\cos(2\phi)\mathbbm{1}-\{B_{0},B_{1}\}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
We obtain the claim by taking the expectation over $\chi:\mathsf{Enc}(x=0)=\chi$, summing over $\alpha$ and applying 3. ∎
To define the isometry, we need to define unitary operators $\tilde{Z}_{B}$ and $\tilde{X}_{B}$ from $Z_{B}$ and $X_{B}$, which maintain the above relations. For this, we follow the regularization procedure of [YN13, MYS12, BP15]: define $Z_{B}^{*}$ from $Z_{B}$ by changing all zero eigenvalues of $Z_{B}$ to 1, and define the operator $\tilde{Z}_{B}:=Z_{B}^{*}|Z_{B}^{*}|^{-1}$. Then $\tilde{Z}_{B}$ is unitary, self-adjoint, commutes with $Z_{B}$ and satisfies $\tilde{Z}_{B}Z_{B}=|Z_{B}|$. We now establish the following:
Claim 5.
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\big{\|}\big{(}\tilde{Z}_{B}-Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{0}$.
Proof.
Using the fact that $\tilde{Z}_{B}^{\dagger}$ is unitary and $\tilde{Z}_{B}Z_{B}=|Z_{B}|$,
$\|\big{(}\tilde{Z}_{B}-Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}=\|\big{(}\mathbbm{1}-|(-1)^{\mathsf{Dec}(\alpha)}Z_{B}|\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
For any self-adjoint operator $A$, we have $\|(\mathbbm{1}-|A|)|\psi\rangle\|^{2}=\langle\psi|(\mathbbm{1}+|A|^{2}-2|A|)|\psi\rangle\leq\langle\psi|(\mathbbm{1}+|A|^{2}-2A)|\psi\rangle=\|(\mathbbm{1}-A)|\psi\rangle\|^{2}$. Applying this with $A=(-1)^{\mathsf{Dec}(\alpha)}Z_{B}$, taking the expectation over $\chi:\mathsf{Enc}(x=0)=\chi$ and summing over $\alpha$, we can directly applying 1 to get the desired bound. ∎
Claim 6.
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\big{\|}\big{(}\tilde{X}_{B}-X_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{1}/\tan^{2}(\phi)=:\delta_{4}$.
Proof.
Note that
$\|\big{(}\tilde{X}_{B}-X_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$ $=\|\big{(}\mathbbm{1}-|X_{B}|\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\leq\|\big{(}\mathbbm{1}+|X_{B}|\big{)}\big{(}\mathbbm{1}-|X_{B}|\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2},$
where we used the same steps in 5 for the first equality, and the operator inequality $\mathbbm{1}+|X_{B}|\geq\mathbbm{1}$ for the inequality. Notice that $|X_{B}|^{2}=(X_{B})^{2}$, which satisfies $\cos^{2}(\phi)(Z_{B})^{2}+\sin^{2}(\phi)(X_{B})^{2}=\mathbbm{1}$ by definition. Plugging this in we find that
$\|\big{(}\mathbbm{1}+|X_{B}|\big{)}\big{(}\mathbbm{1}-|X_{B}|\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$ $=\frac{1}{\tan^{2}(\phi)}\|\big{(}\mathbbm{1}-(Z_{B})^{2}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2},$
and taking the expectation over $\chi:\mathsf{Enc}(x=0)=\chi$, summing over $\alpha$ and applying 2 we obtain the desired bound. ∎
Next, we define the operators $P_{B}^{b}=\frac{1}{2}(\mathbbm{1}+(-1)^{b}\tilde{Z}_{B})$ for $b\in\{0,1\}$ and note a key property.
Claim 7.
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\big{\|}\big{(}P_{B}^{b}-\delta_{\mathsf{Dec}(\alpha),b}\mathbbm{1}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{0}/2$.
Proof.
Note that
$\|\big{(}P_{B}^{b}-\delta_{\mathsf{Dec}(\alpha),b}\mathbbm{1}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$ $=\frac{1}{4}\|\big{(}\tilde{Z}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2},$
where we used the definition of $P_{B}^{b}$ and $\delta_{\mathsf{Dec}(\alpha),b}=\frac{1}{2}(1+(-1)^{\mathsf{Dec}(\alpha)+b})$. Applying the triangle inequality,
$\|\big{(}\tilde{Z}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\|\big{(}Z_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}+\|\big{(}\tilde{Z}_{B}-Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2},$
and taking the expectation over $\chi:\mathsf{Enc}(x=0)=\chi$, summing over $\alpha$ and applying 1 and 5 we arrive at the bound. ∎
Having established the above relations for the case $\chi=\mathsf{Enc}(x=0)$, we are now ready to define the isometry $V:\mathcal{H}\rightarrow\mathbb{C}^{2}\otimes\mathcal{H}$,
$V=\sum_{b\in\{0,1\}}|b\rangle\otimes(\tilde{X}_{B})^{b}P_{B}^{b}.$ (4.14)
This isometry is based on the “SWAP gate isometry” [MYS12, YN13]. The isometry is efficient according to Definition 2.2 and has several applications in the bipartite setting. Using 7, we establish the following lemma:
Lemma 4.5.
Let $\widetilde{\mathrm{Q}}$ be any compiled model which satisfies $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S_{\theta,\phi}]\geq\eta_{\theta,\phi}^{\mathrm{Q}}-\epsilon$, and let $V$ be the isometry defined in Eq. 4.14. Then there exists auxiliary states $\{|\mathsf{aux}_{\alpha|\chi}\rangle\}_{\alpha,\chi}$ such that
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\big{\|}V|\Psi_{\alpha|\chi}\rangle-|\mathsf{Dec}(\alpha)\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{0}.$ (4.15)
Proof.
Consider the states $|\mathsf{aux}_{\alpha|\chi}\rangle=(\tilde{X}_{B})^{\mathsf{Dec}(\alpha)}|\Psi_{\alpha|\chi}\rangle$. Then
$|\mathsf{Dec}(\alpha)\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle$ $=\sum_{b\in\{0,1\}}\delta_{b,\mathsf{Dec}(\alpha)}|b\rangle\otimes(\tilde{X}_{B})^{b}|\Psi_{\alpha|\chi}\rangle,$
and we have
$\big{\|}V|\Psi_{\alpha|\chi}\rangle-|\mathsf{Dec}(\alpha)\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big{\|}^{2}$ $=\big{\|}\sum_{b\in\{0,1\}}|b\rangle\otimes(\tilde{X}_{B})^{b}\big{(}P_{B}^{b}-\delta_{b,\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\leq\sum_{b\in\{0,1\}}\big{\|}\big{(}P_{B}^{b}-\delta_{b,\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
By taking the expectation over $\chi:\mathsf{Enc}(x=0)=\chi$, and summing over $\alpha$, we can apply 7 to obtain the desired bound. ∎
We now proceed to study the relations imposed by the second SOS polynomial, which imply $\hat{E}_{1}\leq\delta_{0}/\lambda_{\theta,\phi}^{2}$. Before we begin, we note that in any earlier claim that did not include a factor of $(-1)^{\mathsf{Dec}(\alpha)}$, we can always apply Lemma 3.1 to exchange the expectation over encryptions of $x=0$ to an expectation over encryptions of $x=1$ at the cost of adding $\delta_{\mathsf{QHE}}$ to the bound. This follows from the fact that such claims are bounding the expectation of operators $\mathcal{N}$ constructed from sums of products of $B_{0}$ and $B_{1}$ only along with the fact that $\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\mathcal{N}|\Psi_{\alpha|\chi}\rangle\approx_{\delta_{\mathsf{QHE}}}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\mathcal{N}|\Psi_{\alpha|\chi}\rangle$.
Claim 8.
$\hat{E}_{1}=\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\big{\|}\big{(}\mathbbm{1}-(-1)^{\mathsf{Dec}(\alpha)}\sin(2\theta)X_{B}-\cos(2\theta)Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{0}/\lambda_{\theta,\phi}^{2}$.
Claim 9.
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\big{\|}\big{(}\mathbbm{1}-(-1)^{\mathsf{Dec}(\alpha)}\sin(2\theta)\tilde{X}_{B}-\cos(2\theta)\tilde{Z}_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{5}$, where $\delta_{5}:=\delta_{0}/\lambda_{\theta,\phi}^{2}+\sin^{2}(2\theta)(\delta_{4}+\delta_{\mathsf{QHE}})+\cos^{2}(2\theta)(\delta_{0}+\delta_{\mathsf{QHE}})$.
Proof.
Using the triangle inequality we obtain
$\big{\|}\big{(}\mathbbm{1}-$ $(-1)^{\mathsf{Dec}(\alpha)}\sin(2\theta)\tilde{X}_{B}-\cos(2\theta)\tilde{Z}_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\leq\big{\|}\big{(}\mathbbm{1}-(-1)^{\mathsf{Dec}(\alpha)}\sin(2\theta)X_{B}-\cos(2\theta)Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\big{\|}\sin(2\theta)\big{(}X_{B}-\tilde{X}_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}+\big{\|}\cos(2\theta)\big{(}Z_{B}-\tilde{Z}_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
Next we take the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$, and sum over $\alpha$. This allows us to apply 8 to the first term, Lemma 3.1 followed by 6 to the second, and Lemma 3.1 followed by 5 to the final term, giving the desired bound. ∎
Claim 10.
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\big{\|}\{\tilde{Z}_{B},\tilde{X}_{B}\}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{6}$, where $\delta_{6}:=\delta_{0}+\delta_{\mathsf{QHE}}+2(\delta_{4}+\delta_{\mathsf{QHE}})+(\delta_{0}+\delta_{\mathsf{QHE}})/(2\sin^{2}(\phi))+(1+1/(2\cos^{2}(\phi)))\delta_{0}/(\sin(2\theta)\lambda_{\theta,\phi})^{2}+(1/\sin^{2}(2\theta)+1/(2\cos^{2}(\phi)\tan^{2}(2\theta)))(\delta_{0}+\delta_{\mathsf{QHE}})$.
Proof.
We begin by applying the triangle inequality to obtain
$\big{\|}\{\tilde{Z}_{B},\tilde{X}_{B}\}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\big{\|}\tilde{X}_{B}\big{(}\tilde{Z}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}+\big{\|}\big{(}\tilde{Z}_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}\tilde{X}_{B}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
For the first term, we can use the fact that $\tilde{X}_{B}$ is unitary, and when taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and sum over $\alpha$, we can apply Lemma 3.1 followed by 1 to obtain a bound of $\delta_{0}+\delta_{\mathsf{QHE}}$. For the second term,
$\big{\|}\big{(}\tilde{Z}_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}\tilde{X}_{B}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$ $\leq\big{\|}\big{(}\tilde{Z}_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}(\tilde{X}_{B}-X_{B})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\|\big{(}\tilde{Z}_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}X_{B}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\leq\big{\|}\tilde{Z}_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\|_{\mathrm{op}}^{2}\big{\|}(\tilde{X}_{B}-X_{B})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\|\big{(}Z_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}X_{B}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\|\big{(}\tilde{Z}_{B}-Z_{B}\big{)}X_{B}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
Since $\tilde{Z}_{B}$ is unitary $\big{\|}\tilde{Z}_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\|_{\mathrm{op}}^{2}\leq 2$, hence when taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and sum over $\alpha$, we can apply Lemma 3.1 followed by 6 to $\|(\tilde{X}_{B}-X_{B})|\Psi_{\alpha|\chi}\rangle\|^{2}$; the product results in a contribution $2(\delta_{4}+\delta_{\mathsf{QHE}})$. Next, we note that
$\big{\|}\big{(}Z_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}X_{B}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$ $=\big{\|}X_{B}\big{(}Z_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\leq\|X_{B}\|_{\mathrm{op}}^{2}\|\big{(}Z_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2},$
where the equality follows since $\{Z_{B},X_{B}\}=0$ by definition. Now $\|X_{B}\|_{\mathrm{op}}^{2}\leq 1/(2\sin^{2}(\phi))$, and when taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and sum over $\alpha$, we can apply Lemma 3.1 followed by 1, resulting in a contribution of $(\delta_{0}+\delta_{\mathsf{QHE}})/(2\sin^{2}(\phi))$.
We now address the term $\|\big{(}\tilde{Z}_{B}-Z_{B}\big{)}X_{B}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$. 8 tells us that
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\big{\|}\big{(}X_{B}-(-1)^{\mathsf{Dec}(\alpha)}T_{\theta,\phi}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{0}/(\sin(2\theta)\lambda_{\theta,\phi})^{2},$ (4.16)
where $T_{\theta,\phi}=\frac{1}{\sin(2\theta)}(\mathbbm{1}-\cos(2\theta)Z_{B})$, which implies
$\|\big{(}\tilde{Z}_{B}-Z_{B}\big{)}X_{B}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$ $\leq\|\big{(}\tilde{Z}_{B}-Z_{B}\big{)}T_{\theta,\phi}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\big{\|}\big{(}\tilde{Z}_{B}-Z_{B}\big{)}\big{(}X_{B}-(-1)^{\mathsf{Dec}(\alpha)}T_{\theta,\phi}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\leq\|\big{(}\tilde{Z}_{B}-Z_{B}\big{)}T_{\theta,\phi}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\big{\|}\tilde{Z}_{B}-Z_{B}\big{\|}_{\mathrm{op}}^{2}\big{\|}\big{(}X_{B}-(-1)^{\mathsf{Dec}(\alpha)}T_{\theta,\phi}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
We have that $\big{\|}\tilde{Z}_{B}-Z_{B}\big{\|}_{\mathrm{op}}^{2}\leq 1+1/(2\cos^{2}(\phi))$, and when taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and summing over $\alpha$, we can apply Eq. 4.16 to obtain a contribution of $(1+1/(2\cos^{2}(\phi)))\delta_{0}/(\sin(2\theta)\lambda_{\theta,\phi})^{2}$. Finally, we bound the last term using the fact that $\tilde{Z}_{B}$ and $Z_{B}$ commute, that is
$\|\big{(}\tilde{Z}_{B}-Z_{B}\big{)}T_{\theta,\phi}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$ $=\big{\|}T_{\theta,\phi}\big{(}\tilde{Z}_{B}-Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\leq\|T_{\theta,\phi}\|^{2}_{\mathrm{op}}\big{\|}\big{(}\tilde{Z}_{B}-Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
The operator norm is bounded by $1/\sin^{2}(2\theta)+1/(2\cos^{2}(\phi)\tan^{2}(2\theta))$, and when taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and summing over $\alpha$, we can apply Lemma 3.1 followed by 5, resulting in an overall contribution of $(1/\sin^{2}(2\theta)+1/(2\cos^{2}(\phi)\tan^{2}(2\theta)))(\delta_{0}+\delta_{\mathsf{QHE}})$. By summing all the contributions from every use of the triangle inequality, we obtain the desired bound. ∎
Claim 11.
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\big{\|}(\tilde{X}_{B}P_{B}^{1}-P_{B}^{0}\tilde{X}_{B})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{6}/4$.
Proof.
Note that
$\tilde{X}_{B}P_{B}^{1}-P_{B}^{0}\tilde{X}_{B}=-\frac{1}{2}\{\tilde{Z}_{B},\tilde{X}_{B}\},$
and hence the claim immediately follows from 10. ∎
We are now ready to prove the self-testing statement on the state when $\mathsf{Enc}(x=1)=\chi$:
Lemma 4.6.
Let $\widetilde{\mathrm{Q}}$ be any compiled model which satisfies $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S_{\theta,\phi}]\geq\eta_{\theta,\phi}^{\mathrm{Q}}-\epsilon$, and let $V$ be the SWAP isometry defined in Eq. 4.14. Then there exists auxiliary states $\{|\mathsf{aux}_{\alpha|\chi}\rangle\}_{\alpha,\chi}$ such that
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\big{\|}V|\Psi_{\alpha|\chi}\rangle-\big{(}\cos(\theta)|0\rangle+(-1)^{\mathsf{Dec}(\alpha)}\sin(\theta)|1\rangle\big{)}\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{7},$ (4.17)
where $\delta_{7}:=\delta_{6}/4+\delta_{5}/(2\sin^{2}(2\theta))$.
Proof.
Consider the choice $|\mathsf{aux}_{\alpha|\chi}\rangle=\frac{1}{\cos(\theta)}P_{B}^{0}|\Psi_{\alpha|\chi}\rangle$. Then
[…]
$=\big{\|}|1\rangle\otimes\big{(}\tilde{X}_{B}P_{B}^{1}-(-1)^{\mathsf{Dec}(\alpha)}\tan(\theta)P_{B}^{0}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\leq\big{\|}\big{(}P_{B}^{0}\tilde{X}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\tan(\theta)P_{B}^{0}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}+\big{\|}(\tilde{X}_{B}P_{B}^{1}-P_{B}^{0}\tilde{X}_{B})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
Taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and summing over $\alpha$, we can bound the second term by $\delta_{6}/4$ using 11. Using the triangle inequality again,
[…]
$\leq\big{\|}\big{(}(-1)^{\mathsf{Dec}(\alpha)}P_{B}^{0}\tilde{T}_{\theta,\phi}-(-1)^{\mathsf{Dec}(\alpha)}\tan(\theta)P_{B}^{0}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\big{\|}P_{B}^{0}\big{(}\tilde{X}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\tilde{T}_{\theta,\phi}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\leq\big{\|}\big{(}P_{B}^{0}\tilde{T}_{\theta,\phi}-\tan(\theta)P_{B}^{0}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\big{\|}P_{B}^{0}\big{\|}_{\mathrm{op}}^{2}\big{\|}\big{(}\tilde{X}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\tilde{T}_{\theta,\phi}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2},$
where $\tilde{T}_{\theta,\phi}=\frac{1}{\sin(2\theta)}(\mathbbm{1}-\cos(2\theta)\tilde{Z}_{B})$. Recall that $\big{\|}P_{B}^{0}\big{\|}_{\mathrm{op}}^{2}\leq 1/2$, and after taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and summing over $\alpha$, we can bound the second term using 9, resulting in a contribution of $\delta_{5}/(2\sin^{2}(2\theta))$ to the final bound. Finally, by direct calculation we obtain
$P_{B}^{0}\tilde{T}_{\theta,\phi}=\frac{1}{\sin(2\theta)}(P_{B}^{0}-\cos(2\theta)P_{B}^{0}\tilde{Z}_{B})=\frac{1-\cos(2\theta)}{{\sin(2\theta)}}P_{B}^{0}=\tan(\theta)P_{B}^{0},$ (4.18)
which implies $\big{\|}\big{(}P_{B}^{0}\tilde{T}_{\theta,\phi}-\tan(\theta)P_{B}^{0}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}=0$. Summing the contributions from each application of the triangle inequality we arrive at the final bound. ∎
Eqs. 4.15 and 4.17 complete the self-testing statement of the model without the action of measurements in the second round according to Eq. 3.21.
Extracting the measurements: Next we turn our attention to the measurements.
Claim 12.
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\big{\|}(VZ_{B}-(\sigma_{Z}\otimes\mathbbm{1})V)|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{8,x}$, where $\delta_{8,x}=\delta_{0}+x\cdot\delta_{\mathsf{QHE}}$ for $x\in\{0,1\}$.
Proof.
We first notice that
$(\sigma_{Z}\otimes\mathbbm{1})V|\Psi_{\alpha|\chi}\rangle$ $=\sum_{b\in\{0,1\}}(-1)^{b}|b\rangle\otimes(\tilde{X}_{B})^{b}P_{B}^{b}|\Psi_{\alpha|\chi}\rangle$
$=\sum_{b\in\{0,1\}}|b\rangle\otimes(\tilde{X}_{B})^{b}P_{B}^{b}\tilde{Z}_{B}|\Psi_{\alpha|\chi}\rangle,$
where we used the fact that $P_{B}^{b}\tilde{Z}_{B}=(-1)^{b}P_{B}^{b}$. Thus,
$\big{\|}(VZ_{B}-$ $(\sigma_{Z}\otimes\mathbbm{1})V)|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}=\big{\|}\sum_{b\in\{0,1\}}|b\rangle\otimes(\tilde{X}_{B})^{b}P_{B}^{b}(Z_{B}-\tilde{Z}_{B})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\leq\Bigg{(}\sum_{b\in\{0,1\}}\|(\tilde{X}_{B})^{b}P_{B}^{b}\|_{\mathrm{op}}^{2}\Bigg{)}\big{\|}(Z_{B}-\tilde{Z}_{B})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\leq\|(Z_{B}-\tilde{Z}_{B})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2},$
For the second, we used the fact that $\|P_{B}^{b}\|_{\mathrm{op}}^{2}\leq 1/2$ and $\|\tilde{X}_{B}P_{B}^{1}\|_{\mathrm{op}}=\|P_{B}^{1}\|_{\mathrm{op}}$ since $\tilde{X}_{B}$ is unitary. For the case $\mathsf{Enc}(x=0)=\chi$, taking the expectation, summing over $\alpha$ and applying 5 recovers the desired bound. For the case $\mathsf{Enc}(x=1)=\chi$, the same procedure with an intermediate application of Lemma 3.1, which accumulates an extra factor of $\delta_{\mathsf{QHE}}$, obtains the desired bound. ∎
Claim 13.
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\big{\|}(VX_{B}-(\sigma_{X}\otimes\mathbbm{1})V)|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{9,x}$, where $\delta_{9,x}=\delta_{4}+\delta_{\mathsf{QHE}}\cdot(2-x)+\delta_{6}/2$.
Proof.
By the triangle inequality
$\big{\|}(VX_{B}-(\sigma_{X}\otimes\mathbbm{1})V)|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$ $\leq\big{\|}V(X_{B}-\tilde{X}_{B})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\big{\|}(V\tilde{X}_{B}-(\sigma_{X}\otimes\mathbbm{1})V)|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
For the first term,
$\big{\|}V(X_{B}-\tilde{X}_{B})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\Bigg{(}\sum_{b\in\{0,1\}}\|(\tilde{X}_{B})^{b}P_{B}^{b}\|_{\mathrm{op}}^{2}\Bigg{)}\big{\|}(X_{B}-\tilde{X}_{B})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
by the same calculation in 12, which after taking the expectation over $\mathsf{Enc}(x)=\chi$ and summing over $\alpha$, results in a contribution of $\delta_{4}$ when $x=0$ by 6, and $\delta_{4}+\delta_{\mathsf{QHE}}$ by Lemma 3.1 and 6 when $x=1$.
For the second term we explicitly calculate
$V\tilde{X}_{B}-(\sigma_{X}\otimes\mathbbm{1})V=|0\rangle\otimes\big{(}P_{B}^{0}\tilde{X}_{B}-\tilde{X}_{B}P_{B}^{1}\big{)}+|1\rangle\otimes\big{(}\tilde{X}_{B}P_{B}^{1}\tilde{X}_{B}-P_{B}^{0}\big{)}.$
Applying the triangle inequality, we end up with the sum of two terms. The first is given by
$\big{\|}(P_{B}^{0}\tilde{X}_{B}-\tilde{X}_{B}P_{B}^{1})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2},$
which after taking the expectation over $\mathsf{Enc}(x)=\chi$ and summing over $\alpha$ results in a contribution of $\delta_{6}/4$ for $x=1$ by 11, and $\delta_{6}/4+\delta_{\mathsf{QHE}}$ for $x=0$ by Lemma 3.1 and 11. By the unitarity of $\tilde{X}_{B}$ the second is equal to $\big{\|}(P_{B}^{1}\tilde{X}_{B}-\tilde{X}_{B}P_{B}^{0})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$. This is bounded by the same quantity as the previous, by noting that the proof of 11 is nearly identical with only the values $b$ of $P_{B}^{b}$ switched. ∎
We are now ready to prove the final self-testing statement.
Lemma 4.7.
Let $\widetilde{\mathrm{Q}}$ be any compiled model which satisfies $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S_{\theta,\phi}]\geq\eta_{\theta,\phi}^{\mathrm{Q}}-\epsilon$, $\widetilde{\mathrm{Q}}^{*}$ be the honest compiled model derived from Eq. 4.5, and let $V$ be the SWAP isometry defined in Eq. 4.14. Then there exists auxiliary states $\{|\mathsf{aux}_{\alpha|\chi}\rangle\}_{\alpha,\chi}$ such that
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\big{\|}VN_{b|y}|\Psi_{\alpha|\chi}\rangle-Q_{b|y}|\varphi_{\mathsf{Dec}(\alpha)|\chi}\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big{\|}^{2}\leq\zeta_{x},$ (4.19)
where $\zeta_{x}=(\cos^{2}(\phi)\,\delta_{8,x}+\sin^{2}(\phi)\,\delta_{9,x})/4+(1-x)\cdot\delta_{7}+x\cdot\delta_{0}$, and $(\mathbb{C}^{2},\{|\varphi_{a|x}\rangle\},\{Q_{b|y}\})$ is the compiled counterpart of the bipartite model defined in Eq. 4.5.
Proof.
By definition $N_{b|y}=(1/2)(\mathbbm{1}+(-1)^{b}[\cos(\phi)\,Z_{B}+(-1)^{y}\sin(\phi)\,X_{B}])$. We also have that the target measurements are given by $Q_{b|y}=(1/2)(\mathbbm{1}+(-1)^{b}[\cos(\phi)\,\sigma_{Z}+(-1)^{y}\sin(\phi)\,\sigma_{X}])$ in Eq. 4.5. We first calculate
$VN_{b|y}|\Psi_{\alpha|\chi}\rangle-(Q_{b|y}\otimes\mathbbm{1})V|\Psi_{\alpha|\chi}\rangle$ $=\frac{(-1)^{b}\cos(\phi)}{2}(VZ_{B}-(\sigma_{Z}\otimes\mathbbm{1})V)|\Psi_{\alpha|\chi}\rangle$
$+\frac{(-1)^{b+y}\sin(\phi)}{2}(VX_{B}-(\sigma_{X}\otimes\mathbbm{1})V)|\Psi_{\alpha|\chi}\rangle.$
Taking the vector norm of both sides, applying the triangle inequality, then taking the expectation over $\mathsf{Enc}(x)=\chi$ and summing over $\alpha$, we arrive at
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\big{\|}(VN_{b|y}-(Q_{b|y}\otimes\mathbbm{1})V)|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\frac{\cos^{2}(\phi)\,\delta_{8,x}+\sin^{2}(\phi)\,\delta_{9,x}}{4}$ (4.20)
by 12 and 13. Finally, we note that Eqs. 4.15 and 4.17 imply the existence of states $\{|\mathsf{aux}_{\alpha|\chi}^{\prime}\rangle\}_{\alpha,\chi}$ such that
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\big{\|}(Q_{b|y}\otimes\mathbbm{1})V|\Psi_{\alpha|\chi}\rangle-Q_{b|y}|\mathsf{Dec}(\alpha)\rangle\otimes|\mathsf{aux}_{\alpha|\chi}^{\prime}\rangle\big{\|}^{2}\leq\delta_{0},$ (4.21)
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\big{\|}(Q_{b|y}\otimes\mathbbm{1})V|\Psi_{\alpha|\chi}\rangle-Q_{b|y}\big{(}\cos(\theta)|0\rangle+(-1)^{\mathsf{Dec}(\alpha)}\sin(\theta)|1\rangle\big{)}\otimes|\mathsf{aux}_{\alpha|\chi}^{\prime}\rangle\big{\|}^{2}\leq\delta_{7},$
where we used that vector norm can only decrease under the application of a projector $Q_{b|y}\otimes\mathbbm{1}$. Note that for the honest strategy in Eq. 4.5, the compiled counterpart consists of the sub-normalized states
$|\varphi_{a|x=0}\rangle$ =$=\begin{cases}\cos(\theta)|0\rangle_{B}\ \mathrm{if}\ a=0,\\ \sin(\theta)|1\rangle_{B}\ \mathrm{if}\ a=1,\end{cases}$
$|\varphi_{a|x=1}\rangle$ $=\frac{\cos(\theta)|0\rangle_{B}+(-1)^{a}\sin(\theta)|1\rangle_{B}}{\sqrt{2}}.$
For $\chi:\mathsf{Enc}(x=0)=\chi$, define $|\mathsf{aux}_{\alpha|\chi}\rangle:=|\mathsf{aux}_{\alpha|\chi}^{\prime}\rangle/\cos(\theta)$ for $\alpha:\mathsf{Dec}(\alpha)=0$, and $|\mathsf{aux}_{\alpha|\chi}\rangle:=|\mathsf{aux}_{\alpha|\chi}^{\prime}\rangle/\sin(\theta)$ for $\alpha:\mathsf{Dec}(\alpha)=1$. For $\chi:\mathsf{Enc}(x=1)=\chi$, define $|\mathsf{aux}_{\alpha|\chi}\rangle:=\sqrt{2}|\mathsf{aux}_{\alpha|\chi}^{\prime}\rangle$ for all $\alpha$. Then the triangle inequality implies
$\big{\|}VN_{b|y}|\Psi_{\alpha|\chi}\rangle-$ $Q_{b|y}|\varphi_{\mathsf{Dec}(\alpha)|x}\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big{\|}^{2}\leq\big{\|}VN_{b|y}|\Psi_{\alpha|\chi}\rangle-(Q_{b|y}\otimes\mathbbm{1})V|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\big{\|}(Q_{b|y}\otimes\mathbbm{1})V|\Psi_{\alpha|\chi}\rangle-Q_{b|y}|\varphi_{\mathsf{Dec}(\alpha)|\chi}\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big{\|}^{2},$
which is bounded by Eqs. 4.20 and 4.21, after taking the expectation over $\mathsf{Enc}(x)=\chi$ and summing over $\alpha$, by $\zeta_{x}=(\cos^{2}(\phi)\,\delta_{8,x}+\sin^{2}(\phi)\,\delta_{9,x})/2+(1-x)\cdot\delta_{7}+x\cdot\delta_{0}$. ∎
Proof of Theorem 4.4..
Eq. 4.19 shows that the generalized tilted-CHSH expressions are a self-test according to Eq. 3.21, completing the proof of Theorem 4.4. ∎